### v2

# MDS-ED: Multimodal Decision Support

Abstract
Background: Benchmarking medical decision support algorithms often struggles due to limited access to datasets, narrow prediction tasks, and restricted input modalities. These limitations affect their clinical relevance and performance in high-stakes areas like emergency care, complicating replication, validation, and improvement of benchmarks.
Methods: We introduce a dataset based on MIMIC-IV, benchmarking protocol, and initial results for evaluating multimodal decision support in the emergency department (ED). We use diverse data modalities from the first 1.5 hours of patient arrival, including demographics, biometrics, vital signs, lab values, and electrocardiogram waveforms. We analyze 1443 clinical labels across two contexts: predicting diagnoses with ICD-10 codes and forecasting patient deterioration.
Results: Our multimodal diagnostic model achieves an AUROC score over 0.8 in a statistically significant manner for 357 out of 1428 conditions, including cardiac issues like myocardial infarction and non-cardiac conditions such as renal disease and diabetes. The deterioration model scores above 0.8 in a statistically significant manner for 13 out of 15 targets, including critical events like cardiac arrest and mechanical ventilation, ICU admission as well as short- and long-term mortality. Incorporating raw waveform data significantly improves model performance, which represents one of the first robust demonstrations of this effect.
Conclusions: This study highlights the uniqueness of our dataset, which encompasses a wide range of clinical tasks and utilizes a comprehensive set of features collected early during the emergency after arriving at the ED. The strong performance, as evidenced by high AUROC scores across diagnostic and deterioration targets, underscores the potential of our approach to revolutionize decision-making in acute and emergency medicine.
1 Introduction
Artificial intelligence (AI) is an emerging field that has significantly enhanced healthcare and medicine [1] in areas such as precision medicine [2] and drug discovery [3]. The implementation of AI models is particularly relevant in acute and emergency medicine, where clinicians must address critical conditions within short time frames to make optimal clinical decisions. AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8]. There has been exponential growth in the number of scientific publications on new AI models. However, many of these studies have limitations, such as predicting only a limited set of specific conditions [9, 4], offering short time horizons for survival predictions, and requiring costly diagnostic tests [10].
Healthcare datasets are essential for advancing medical research and innovation. However, much of this valuable data remains undisclosed due to commercial interests and privacy concerns. Despite these hurdles, there has been a recent surge in publicly accessible datasets aimed at facilitating the development and validation of machine learning models to tackle complex healthcare challenges [11, 12, 13, 14, 15]. Nonetheless, these datasets still have limitations, including size constraints [12], lack of open-source availability [16, 17], and narrow scopes in terms of prediction tasks [12, 13, 17, 14, 16, 15]. Benchmarking is crucial for evaluating the performance and reliability of novel AI-models, ensuring their applicability and robustness in real-world clinical settings. In this study, we investigate multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
2 Methods
2.1 Clinical workflow and dataset creation
Figure 1: Pipeline outlines the MDS-ED workflow, which involves feature collection encompassing patient demographics, biometrics (such as height, weight, and BMI), vital parameters and trends, laboratory values and trends, and ECG waveform data to address two clinically relevant prediction scenarios: predicting patient discharge diagnoses out of 1428 cardiac and non-cardiac ICD10-CM codes and predicting patient deterioration according to 15 clinical deterioration measures.
Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay. MDS-ED was created by linking ECG waveforms from the MIMIC-IV-ECG [18] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [19].
Prediction tasks and targets
In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration. For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [11], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG). To this end, we frame the task as a multilabel classification task in terms of ICD-10 CM codes. As in [11], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits and propagate codes to parent codes up to the third digit. In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions. For patient deterioration, we build on a consensus definition [20] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA). Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days. With these two main groups, we aim to cover a wide range spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels, see Table 5 in the supplementary material for a detailed target definition discussion.
Features
While aiming to include mainly data at triage, we also include irregularly sampled vital parameters and lab values captured during the 1.5h data collection window. To capture trends, we compute simple sample-wise statistical aggregation functions per vital parameter and laboratory values, namely mean, median, minimum, maximum, standard deviation, first and last values, rate of change of them between first and last value, as well as the slope of a linear model fitted on the minutes difference between value collection and arrival as independent variable and the actual values as dependent variables. We follow [21] to select an appropriate set of laboratory values that avoids a bias towards particular prediction targets. We obtain a total of 470 features across all feature modalities in a tabular format plus a 12-lead 10-second ECG waveform per sample.
MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,095 samples. We leverage the stratified splits provided along with the dataset accompanying [11], which includes stratification based on gender, age bins, and discharge diagnoses. We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1, see Appendix A in the supplementary material for further preprocessing and definitions details as well as descriptive statistics on demographics, biometrics, vital parameters and laboratory values in Tables 6-8.
2.2 Related work
Group Detail MIMIC-ED-Assist VitalML MEME EHRSHOT MC-BEC MDS-ED
Source [12] [13] [17] [14] [16] This work
Population ED ED Longitudinal Longitudinal ED ED
Size Patients 25714 N/A 947028 6739 63389 71098
Visits 32356 19847 N/A 921499 102731 121195
Features Demographics
Biometrics
Vital parameters ✓($\mathbb{T}$ ) ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$)
Lab. values ✓($\mathbb{T}$) ✓($\mathbb{T}$)
Waveforms $(\mathbb{E}$)
Chief complaint
Medications
Tasks Labels 3 3 4 15 7 1443
Availability Open source
Table 1: Direct comparison with related works in terms of dataset size, features, availability, and number of target labels. We use diverse symbology to express the contribution where ✓ = available, ✗ = unavailable, $\mathbb{E}$ = available in the form of embeddings, and $\mathbb{T}$ = available in the form of trends or at least two sampled values.
Table 1 contains a direct comparison with related works in terms of population, dataset size, features, availability, and number of target labels. [12] proposed the MIMIC-ED-Assist benchmark to reduce ED length of stay by flagging high-risk patients using triage features. [13] proposed VitalML to predict deterioration in the next 90 minutes from the first 15 minutes of monitoring. [17] proposed a multiple embedding model for EHR (MEME) to generate pseudo-notes from tabular data and predict ED disposition, discharge location, ICU admission, and inpatient mortality utilizing a natural language processing approach. [14] proposed a few shot foundational model EHRSHOT approach to predict operational outcomes, developed diagnoses, as well as anticipate lab test results and image findings. [16] proposed MC-BEC to predict patient deterioration, disposition, and revisit from features collected including diagnoses from the first 15 minutes after the patient is assigned to a room.
Overall, our proposed MDS-ED main contributions in terms of the dataset are: 1) Comprehensive size: MDS-ED is situated in the first position regarding a large number of patients and second in the number of visits in the open-source domain, despite restricting towards the investigated setting of the first 1.5 hours of ED arrival. 2) Diverse input features: MDS-ED is situated in the first position in terms of feature modalities in the open-source domain as it incorporates a comprehensive set of features, including demographics, biometrics, vital parameter trends, laboratory values trends, and ECG waveforms. This is more extensive than most compared datasets. Intentionally, we decided to exclude chief complaints as their unstructured nature presents a more challenging step towards external validation with the addition of more variance, see [17]. Similarly, we exclude previous patient medications as they involve irregularities in administration times, and dosages, as well as potential bias towards task prediction, e.g., certain medications are given for specific gravity and stages of diagnoses. 3) Comprehensive prediction targets: MDS-ED proposes 1,443 target labels, significantly more than other datasets, which typically cover fewer and too narrow in scope tasks. 4) Accessibility: Similar to [12, 13, 14], MDS-ED is open source, encouraging further research and collaboration.
2.3 Model architectures
To propose an initial benchmark, we experiment with diverse settings based on data modalities input such as (1) tabular only, (2) ECG waveforms only, (3) ECG features + tabular, and (4) ECG waveforms + tabular. For waveforms (2) and (4), inspired by recent successful applications in the field of physiological time series [11, 22, 23] we employed an S4 classifier based on structured state space models with four layers [24], specifically, for the multimodal approach in (4) we extended the S4 classifier by fusing it with the tabular features using a tensor fusion approach based on a differentiable outer product proposed in the context of multimodal sentiment analysis [25]. For (1) and (3) we fit gradient-boosting decision trees (XGBoost). Additional details on model architectures and model hyperparameters can be found in Appendix B Table 9.
2.4 Training and evaluation
For the S4 models, we used AdamW as optimizer, with both learning rate and weight decay set to 0.001, and maintained a constant learning rate schedule. The training was conducted with a batch size of 64 samples over 20 epochs which usually converged earlier and model selection on the validation set. The training objective was to minimize binary cross-entropy loss. For all the models, we evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC). To assess statistical uncertainty resulting from the finite size and specific composition of the test set, we employ empirical bootstrap on the test set with $n=1000$ iterations. We report 95% confidence intervals for both macro AUROC and individual label AUROCs. Given [11] evaluation protocol, we use only the first record per visit per patient to avoid bias. Additional details of training and evaluation procedures can be found in Appendix B.
3 Results
3.1 Benchmarking predictive performance
Diagnoses Deterioration
Tabular only 0.7574 (0.7561, 0.7610) 0.8519 (0.8502, 0.8535)
ECG waveforms only 0.7631 (0.7590, 0.7673) 0.8279 (0.8047, 0.8511)
ECG features + Tabular 0.7684 (0.7673, 0.7721) 0.8693 (0.8688, 0.8715)
ECG waveforms + Tabular 0.7873 (0.7832, 0.7913) 0.8815 (0.8625, 0.8977)
Table 2: Predictive performance (macro AUROC with 95% bootstrap confidence intervals) across all tasks for unimodal models as well as multimodal prediction models.
Table 2 presents the predictive performance of both unimodal and multimodal models across the diagnoses and deterioration settings. For diagnosis prediction, the multimodal model that integrates ECG waveforms and tabular data achieved an AUROC of 0.7873 (0.7832, 0.7913), which significantly outperformed the unimodal models as well as the ECG features and tabular. Interestingly, the unimodal ECG waveforms model outperformed the unimodal tabular model with 0.7631 (0.7590, 0.7673) against 0.7574 (0.7561, 0.7610). In the deterioration prediction task, the multimodal model achieved an AUROC of 0.8815 (0.8625, 0.8977), which is notably higher than the unimodal models as well as the ECG features and tabular. Interestingly, contrary to the diagnoses setting, the tabular-only model outperformed the ECG waveform-only model with 0.8519 (0.8502, 0.8535) against 0.8279 (0.8047, 0.8511). In the following discussion, we restrict ourselves to the discussion of the most comprehensive multimodal model based on tabular features and ECG waveforms.
3.2 Task-dependent predictive performance
Code: AUROC [instances]. Description Code: AUROC [instances]. Description
IX: Circulatory System. 0.8467 (93/181 >0.80 lower bound AUROC)
I2119: 0.9975 [395] ST elevation (STEMI) myocardial infarction involving other coronary artery of inferior wall I132: 0.9803 [871] Hypertensive heart and chronic kidney disease with heart failure and with stage 5 chronic kidney disease, or end stage renal disease
I081: 0.9793 [397] Rheumatic disorders of both mitral and tricuspid valves I447: 0.9691 [1088] Left bundle-branch block, unspecified
I255: 0.9658 [946] Ischemic cardiomyopathy I420: 0.9616 [315] Dilated cardiomyopathy
I481: 0.9549 [265] Persistent atrial fibrillation I120: 0.9525 [3350] Hypertensive chronic kidney disease with stage 5 chronic kidney disease or end-stage renal disease
I30: 0.9514 [316] Acute pericarditis I5023: 0.9494 [3216] Acute on chronic systolic (congestive) heart failure
III: Blood and immune mechanism: 0.8011 (14/45 >0.80 lower bound AUROC)
D6181: 0.9333 [1503] Pancytopenia D631: 0.9268 [2960] Anemia in chronic kidney disease
D6832: 0.9194 [271] Hemorrhagic disorder due to extrinsic circulating anticoagulants D69: 0.8712 [5599] Purpura and other hemorrhagic conditions
IV: Endocrine, nutritional and metabolic. 0.801 (38/107 >0.80 lower bound AUROC)
E1143: 0.9529 [249] Type 2 diabetes mellitus with diabetic autonomic (poly)neuropathy E1022: 0.9393 [309] Type 1 diabetes mellitus with diabetic chronic kidney disease
E43: 0.9216 [1417] Unspecified severe protein-calorie malnutrition E874: 0.9187 [632] Mixed disorder of acid-base balance
XIV: Genitourinary system 0.7978 (12/39 >0.80 lower bound AUROC)
N186: 0.9594 [4362] End stage renal disease N08: 0.9334 [798] Glomerular disorders in diseases classified elsewhere
N170: 0.906 [1773] Acute kidney failure with tubular necrosis N2581: 0.8816 [472] Secondary hyperparathyroidism of renal origin
XIX: Injury, poisoning: 0.797 (20/84 >0.80 lower bound AUROC)
S0180: 0.9143 [261] Unspecified open wound of other part of head S063: 0.9125 [249] Focal traumatic brain injury
T8612: 0.9075 [559] Kidney transplant failure T4551: 0.9021 [572] Poisoning by, adverse effect of and underdosing of anticoagulants
II: Neoplasms. 0.792 (6/53 >0.80 lower bound AUROC)
C228: 0.9326 [236] Malignant neoplasm of liver, primary, unspecified as to type C786: 0.9198 [490] Secondary malignant neoplasm of retroperitoneum and peritoneum
C83: 0.8825 [370] Non-follicular lymphoma C7952: 0.8794 [629] Secondary malignant neoplasm of bone marrow
I: Infectious and parasitic diseases 0.7908 (7/51 >0.80 lower bound AUROC)
A4189 0.9688 [267] Other specified sepsis B9681 0.8749 [245] Helicobacter pylori [H. pylori] as the cause of diseases
X: Respiratory system. 0.7886 (23/71 >0.80 lower bound AUROC)
J9691 0.9530 [572] Respiratory failure, unspecified with hypoxia J152 0.9438 [289] Pneumonia due to staphylococcus
J841 0.9136 [353] Other interstitial pulmonary diseases with fibrosis J948 0.9048 [369] Other specified pleural conditions
XXI: Health status and services 0.7884 (40/163 >0.80 lower bound AUROC)
Z992 0.9681 [3224] Dependence on renal dialysis Z902 0.9633 [265] Acquired absence of lung [part of]
Z681 0.9535 [853] Body mass index (BMI) 19.9 or less, adult Z4502 0.9533 [345] Encounter for adjustment and management of automatic implantable cardiac defibrillator
XI: Digestive system 0.7825 (29/119 >0.80 lower bound AUROC)
K7031: 0.9817 [485] Alcoholic cirrhosis of liver with ascites K7290: 0.9605 [548] Hepatic failure, unspecified without coma
K758: 0.9299 [457] Other specified inflammatory liver diseases K762: 0.9272 [350] Central hemorrhagic necrosis of liver
XII: Skin and subcutaneous tissue 0.7804 (5/25 >0.80 lower bound AUROC)
L9750: 0.9303 [297] Non-pressure chronic ulcer of foot L409: 0.8694 [239] Psoriasis
V: Mental and behavioral disorders 0.7719 (11/71 >0.80 lower bound AUROC)
F500: 0.9856 [262] Anorexia nervosa F1012: 0.9243 [625] Alcohol abuse with intoxication
F4310: 0.9050 [748] Post-traumatic stress disorder, unspecified F1210: 0.8753 [396] Cannabis abuse, uncomplicated
XVIII: Symptoms, signs, and abnormal findings 0.7477 (19 /167 >0.80 lower bound AUROC)
R6521: 0.9561 [2545] Severe sepsis with septic shock R5081: 0.9532 [406] Fever presenting with conditions classified elsewhere
R570: 0.9507 [981] Cardiogenic shock R297: 0.9273 [290] National Institutes of Health Stroke Scale (NIHSS) score
XIII: Musculoskeletal system 0.7282 (8/66 >0.80 lower bound AUROC)
M6282: 0.9224 [598] Rhabdomyolysis M7966: 0.8944 [231] Pain in lower leg
M8580: 0.8931 [540] Other specified disorders of bone density and structure M86 0.8838 [489] Osteomyelitis
Table 3: Best-performing individual statements organized according to selected ICD chapters underscoring the breadth of accurately predictable statements. The table shows the four best-performing individual statements per ICD chapter (10 for chapter IX (Circulatory system diseases), 4 for the rest), where we show only AUROC scores above 0.85 and where also the lower bound of the 95% bootstrap confidence interval exceeds 0.80. To showcase the breadth of reliably predictable statements, we list only the best-performing statement per 3-digit ICD code.
Table 3 demonstrates the diagnoses that can be predicted most accurately by the model, which we organized based on ICD chapters. The algorithm’s predictive performance ranges from an AUROC of 0.7282 for the Musculoskeletal System (XIII) chapter to 0.8467 for the Circulatory System (IX) chapter. In total, the model predicted 325 out of 1,242 individual ICD diagnoses with high accuracy, which we define in this work as conditions where the lower bound of the 95% bootstrap confidence interval for AUROC exceeding 0.80.
Label: AUROC [instances] Label: AUROC [instances]
Clinical deterioration. 0.8620 (4/6 >0.80 lower bound AUROC)
Severe hypoxemia: 0.5839 [555] ECMO: 0.8705 [166]
Vasopressors: 0.9132 [1172] Inotropes: 0.9079 [376]
Mechanical ventilation: 0.9394 [4230] IHCA: 0.9576 [623]
ICU admission. 0.885 (2/2 >0.80 lower bound AUROC)
24-hours: 0.8923 [15688] Overall: 0.8777 [18872]
Mortality. 0.8970 (7/7 >0.80 lower bound AUROC)
In-hospital: 0.8954 [99] 24-hours: 0.9512 [817]
7-days: 0.9149 [2275] 28-days: 0.9009 [5153]
90-days: 0.8850 [9300] 180-days: 0.8692 [12844]
365-days: 0.8629 [17761]
Table 4: Detailed breakdown of model performance in the deterioration category.
Table 4 showcases the deterioration model’s predictive performance across clinical deterioration, ICU admission, as well as both short and long-term mortality with many AUROC scores significantly exceeding the 0.80 threshold. Overall, for clinical deterioration, the model achieves an AUROC of 0.8692, whereas for ICU admissions, the model reports an overall AUROC of 0.8842, and for mortality predictions also exhibits high performance, with an overall AUROC of 0.9030.
4 Discussion
4.1 Impact of data modalities
From Section 3.1 we can draw several conclusions: Firstly, the results demonstrate that multimodal models, which integrate diverse data types, offer superior performance in both diagnostic and deterioration tasks (row 1/2 vs. row 3/4). Secondly, in the diagnoses task, the use of ECG raw waveforms instead of ECG features improves the performance in a statistically significant manner (row 3 vs. row 4), whereas for the deterioration task, we performed a direct comparison via bootstrapping the score difference for statistical significance, and we did not find a statistically significant difference, which is in line with [13]. To the best of our knowledge, this is the first statistically robust demonstration of the added value of raw waveform input against features for clinically relevant prediction tasks such as diagnoses prediction. Thirdly, the model building on ECG waveforms as only input outperforms the tabular-only model in the diagnostic setting, but not in the deterioration setting (row 1 vs. row 2). We hypothesize that this is due to the inclusion of tabular trends over time which aligns with the task definition of deterioration. We believe that the inclusion of multiple raw ECGs over time instead of just a single snapshot would allow us to capture more meaningful deterioration trends also from raw waveform data.
4.2 Clinical significance
From Section 3.2, we conclude that our diagnostic models exhibit high predictive capabilities for both acute and chronic conditions. Acute conditions like ST elevation myocardial infarction, acute pericarditis, persistent atrial fibrillation, and acute kidney failure require rapid diagnosis and immediate intervention to prevent severe outcomes. The model’s strong performance in identifying these conditions highlights its potential to significantly impact patient care through timely and accurate diagnosis. Conversely, chronic conditions such as hypertensive heart and chronic kidney disease, ischemic cardiomyopathy, dilated cardiomyopathy, and end-stage renal disease benefit from ongoing management and long-term treatment strategies. The model’s ability to effectively predict these chronic conditions underscores its utility in supporting sustained patient monitoring and management, ultimately improving patient outcomes through consistent and reliable diagnostic support. As demonstrated in [11], even the model based only on ECG waveforms can reliably predict cardiac but most notably non-cardiac conditions. This applies even more to the multimodal modal that involves clinical features as well as ECG waveforms.
Our deterioration model has major clinical relevance by enhancing early recognition of patient deterioration and supporting physician decisions in the emergency department. Its high predictive performance across various scenarios effectively identifies patients at risk for severe complications, which is crucial in acute and emergency medicine. The model’s ability to reliably predict clinically relevant indicators, such as vasopressor use, inotrope use, mechanical ventilation, and in-hospital cardiac arrest, enables healthcare providers to anticipate these events and implement early, tailored interventions. Additionally, its robust performance in predicting ICU admissions allows for better resource allocation, ensuring that patients who need intensive care receive it promptly while preventing unnecessary admissions. The model accurately predicts 24-hour in-hospital mortality and has high accuracy in predicting 365-day mortality, supporting clinical decision-making in prioritizing high-risk patients and discussing prognosis with families. Long-term mortality predictions aid in advanced care planning, such as deciding against ICU admission for patients with a short expected lifespan after discharge, and inform post-discharge planning and follow-up care to prevent readmissions and improve long-term health outcomes.
4.3 Limitations
Despite the promising results, there are several limitations to consider. Firstly, the presented models rely on the specifics of Beth Israel Deaconess Medical Center in terms of patient demographics, available resources, process characteristics, and data collection practices. We carefully quantify the statistical uncertainties of our results and are convinced of the robustness of our results for this particular scenario within the limits of a retrospective evaluation. However, we cannot make any statements about the transferability of our findings to other hospitals or deployment scenarios as this would require an external validation dataset, see [26] for a recent perspective on the role of external validation for clinical decision support. Secondly, the reliance on data collected within the first 1.5 hours of the ED visits may limit the model’s applicability. We deliberately decided to focus on this setting to keep the dataset homogeneous. However, there is also a clear clinical case for a prediction model only incorporating information available during triage. A similar dataset to the one proposed could be created, but this task is left for future research. Thirdly, while we focus on the construction of a benchmark dataset and provide the first results for it, we do not provide any insights into the trained models beyond a discussion of the model performance for specific conditions. This could be achieved for example through attribution methods proposed in the context of explainable AI, however, mostly in uni-modular settings [27], to verify that the models do not exploit spurious correlations to reach their strong performance. Again, a detailed investigation also of causal vs. associational attributions along the lines of [28] is beyond the scope of this submission. Finally, although the open-source dataset fosters innovation, the implementation of a large-scale single end-to-end model with a wide range of target variables such as the diagnoses case might result in challenges during deployment in resource-constrained hardware, see Appendix B in the supplementary material for a more extensive discussion.
4.4 Future research directions
The proposed dataset and benchmark protocol build on input parameters that are typically available in clinical workflows and address a set of comprehensive, clinically meaningful prediction tasks. We therefore envision it to represent a meaningful baseline to benchmark prediction algorithms in the field. An obvious extension would be to include free-text data along the lines of [15]. Going beyond MIMIC-IV, models developed on the proposed dataset could foreshadow the development of more comprehensive models on more specialized datasets such as [29], for example including molecular data as input and/or prediction targets, to address more fine-grained diagnostic or deterioration prediction targets.
4.5 Social impact
Clinically, these models enhance decision-making by providing timely and accurate predictions, reducing diagnostic errors, and improving patient care, especially in high-stakes emergency settings. The open-source nature of the curated dataset fosters collaboration and innovation from diverse research communities, allowing for the development of new models and tools that aim to improve benchmark scores and enable more direct comparisons across models. However, questions of algorithmic fairness remain to be investigated in detail in future work.
4.6 Authors contribution statement
JMLA and NS conceived and designed the project. JMLA conducted the full experimental analyses, with NS supervising them, and HB providing critical revision for clinical intellectual content. JMLA produced the first draft, NS and HB revised it. All authors critically revised the content and approved the final version for publication.
4.7 Code and data availability
Code for experiment reproducibility and dataset instructions in a dedicated code repository [30]. For easier usage, the dataset is currently prepared for submission as a Physionet dataset.
Appendix A Dataset details
Name Labels Details
Diagnoses model (1428 labels, based on [11])
Diagnoses 1428 Discharge diagnoses ICD10-CM codes
Deterioration model (15 labels, based on [20])
Mortalities 6 1, 7, 30, 90, 180, 365 days and in-stay
ICU Admission 2 24 hours and in-stay
Vasopressors 1 Within 24 hours: epinephrine, norepinephrine, vasopressin, dobutamine, dopamine, or phenylephrine
Inotropes 1 Within 24 hours: epinephrine, dobutamine, or dopamine
Mechanical Ventilation 1 Same day or next day ICD9/ICD10 codes: 9670, 9671, 9672, 5A1935Z, 5A1945Z, and 5A1955Z
In-Hospital Cardiac Arrest (IHCA) 1 Same day or next day ICD9/ICD10 codes: I469, 4275, I462, V1253, I468
Extracorporeal Membrane Oxygenation (ECMO) 1 Same day or next day ICD9/ICD10 codes: 3961, 3965, 3966, 5A1221Z, 5A1522G, 5A1522H, 5A15223, 5A1522F, 5A15A2F, 5A15A2G, and 5A15A2H
Severe Hypoxemia 1 Oxygen saturation $\leq$ 85% within 24 hours
Table 5: Definitions of labels for the diagnoses and deterioration models, including 1428 discharge diagnoses based on ICD10-CM codes in the diagnoses model, and 15 labels covering clinical outcomes such as mortalities, ICU admission, vasopressors, inotropes, mechanical ventilation, in-hospital cardiac arrest (IHCA), extracorporeal membrane oxygenation (ECMO), and severe hypoxemia in the deterioration model.
Table 5 contains the considered tasks by description, number of labels, and additional details such as target criteria definition. Apart from diagnoses, we use the rest of the tasks defined as patient deterioration based on previous work from the literature that defines physiological deterioration [20]. Although we consider diverse scores such as SOFA significant, calculating these markers requires a more detailed definition of downstream task prediction, which goes beyond the scope of maintaining homogeneity across tasks in this study. The curation of the dataset starts from MIMIC-IV-ECG-ICD [11], from which we obtained the ICD10 diagnoses codes for the complete MIMIC-IV-ECG [18] dataset, therefore, based on ECG records with available ED stay_id, we select the ones that happened between the patient in time and the 1.5 hours window. Mortalities at finite horizons were computed based on the date of death from [18], and for the computation of mortality per stay we include discharge date times from the admissions table in [19] MIMIC-IV to account for final discharge, all of the mortalities dates were previously validated from the source with a follow-up up to 1 year after the patient’s last recorded discharge. ICU admission in 24 hours and per stay were computed based on the ICU stays table from [19]. Mechanical ventilation and ECMO were computed given the procedures table from MIMIC-IV-ED given the patient admission and procedure date. Cardiac arrest was computed from the diagnosis tables from MIMIC-IV-ED and MIMIC-IV, where we first validated if the patient was discharged with the diagnosis code within the ED or hospital otherwise. We clarify that although mechanical ventilation, ECMO, and cardiac arrest do not provide an exact time of events, we carefully approximate the times with dates given the same day or the day after. Vasopressors and inotropes were computed from the pyxis table from MIMIC-IV-ED given the exact time of the applied medication. Finally, severe hypoxemia was computed from the vital signs table from MIMIC-IV-ED. In the deterioration setting, to not discard samples where the events happened before the 1.5 hours for a single task, e.g. severe hypoxemia within 1.5 hours, we include a special token for the task in the label space which we mask during training and evaluation in order not to negatively affect training and evaluation for other labels due to exclusion of entire patients, see Appendix B for more details.
Name Values
Gender (%)
Female 68289 (52.9)
Male 60806 (47.1)
Age (%)
Median years (SD) 64 (28)
Quantile 1: 18-49 32634 (25.27)
Quantile 2: 50-64 34376 (26.62)
Quantile 3: 65-77 31821 (24.64)
Quantile 4: 78-101 30264 (23.44)
Ethnicity (%)
White 78310 (60.66)
Black 27680 (21.44)
Hispanic 9230 (7.14)
Asian 4700 (3.64)
Other 9175 (7.10)
Table 6: Summary demographic characteristics across samples including gender distribution, age distribution (median with standard deviation and quartiles), and ethnicity distribution in the study population, presented as percentages.
Table 6 provides detailed demographic information on the study population, including gender, age, and ethnicity. The gender distribution shows a slightly higher proportion of females (52.9%) compared to males (47.1%). The age distribution is divided into four quantiles, with a median age of 64 years and a standard deviation of 28 years. Regarding ethnicity, the majority of the population is White (60.66%), followed by Black (21.44%), Hispanic (7.14%), Asian (3.64%), and individuals of other ethnicities (7.10%). This diverse demographic distribution ensures a comprehensive evaluation of the model across various patient groups. These variables were extracted from the ed stays table from MIMIC-IV-ED.
Name Unit Minimum Maximum Median IQR Records
Biometrics
BMI kg/m2 0 99.2 27.7 8.70 70321
Weight kg 20.41 367.99 77.11 27.56 73646
Height cm 64.99 387.09 167.64 15.23 32573
Vital parameters
temperature C 13.3 44.72 36.6 0.55 77179
heartrate bpm 0.0 570.0 84.0 30.0 153249
resprate bpm 0.0 169.0 18.0 4.0 150525
o2sat % 0.0 100.0 99.0 3.0 146140
sbp mmHg 0.0 274.0 130.0 35.0 151346
dbp mmHg 0.0 486.0 73.0 23.0 151254
Triage
Acuity level 1 5 2 1 126071
Table 7: Summary of statistics across samples for biometric measurements, vital parameters, and triage acuity levels, including minimum, maximum, median, interquartile range (IQR), and number of records for each variable.
Table 7 provides details on the biometrics, vital parameters, and triage values considered in this work, including the name, unit, minimum value, maximum value, median value, and the number of records. Dataset preprocessing includes weight and height conversion from pounds and inches to kilograms and centimeters. During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence. The specific thresholds for outliers were as follows: temperature below and above 50 and 150 Fahrenheit degrees respectively. Heart rate above 700, respiration rate above 300, oxygen saturation below 0 and above 100, diastolic and systolic blood pressures above 500. BMI above 100, weight below 20 or above 400, and height below 60 and above 400. The temperature unit was converted from Fahrenheit to Celsius, whereas weight from pounds to kilograms, and height from inches to centimeters. Biometrics features were extracted from the online medical records table from MIMIC-IV from which we match the closest value of a patient stay within 30 days of admission, otherwise missing. Vital parameters and acuity were extracted from the vital signs and triage tables of MIMIC-IV-ED.
Name Unit Minimum Maximum Median IQR Fluid Records
Abs. Basophil Count K/uL 0.0 9.2 0.04 0.03 Blood 36837
Abs. Eosinophil Count K/uL 0.0 15.33 0.09 0.15 Blood 36838
Abs. Lymphocyte Count K/uL 0.0 88.08 1.47 1.12 Blood 36909
Alanine Aminotransf. IU/L 1.0 1999.0 22.0 22.0 Blood 29999
Albumin g/dL 1.0 6.7 4.0 0.79 Blood 28324
Alkaline Phosphatase IU/L 5.0 1985.0 85.0 51.0 Blood 29923
Asparate Aminotransf. IU/L 1.0 1996.0 29.0 29.0 Blood 29995
Bands % 0.0 61.0 0.0 3.0 Blood 6057
Base Excess mEq/L -413.0 29.0 0.0 6.0 Blood 8011
Basophils % 0.0 35.0 0.4 0.39 Blood 79386
Bicarbonate mEq/L 2.0 50.0 24.0 5.0 Blood 79771
Bilirubin, Direct mg/dL 0.0 45.5 1.0 2.5 Blood 1174
Bilirubin, Total mg/dL 0.0 55.8 0.5 0.5 Blood 30098
C-Reactive Protein mg/L 0.1 299.3 17.95 75.14 Blood 1108
Calcium, Total mg/dL 0.0 79.2 9.2 0.79 Blood 30262
Carboxyhemoglobin % 0.0 18.0 2.0 2.0 Blood 1078
Chloride mEq/L 58.0 140.0 102.0 6.0 Blood 79855
Creatine Kinase (CK) IU/L 8.0 1998.0 115.0 158.0 Blood 9588
CK, MB Isoenzyme ng/mL 1.0 497.0 3.0 4.0 Blood 9272
Creatinine mg/dL 0.0 43.0 1.0 0.5 Blood 82850
Eosinophils % 0.0 75.0 1.3 2.1 Blood 79387
Fibrinogen, Functional mg/dL 35.0 1348.0 298.5 161.5 Blood 2280
Free Calcium mmol/L 0.48 3.31 1.08 0.12 Blood 1472
Glucose mg/dL 3.8 1684.0 113.0 50.0 Blood 85022
Hematocrit % 6.9 66.1 38.3 7.89 Blood 82490
Hemoglobin g/dL 0.0 21.8 12.6 3.0 Blood 84890
INR(PT) nan 0.75 24.0 1.1 0.30 Blood 44656
Lactate mmol/L 0.3 25.2 1.8 1.20 Blood 32666
Lymphocytes % 0.0 100.0 19.7 17.1 Blood 79387
Magnesium mg/dL 0.2 45.0 2.0 0.40 Blood 30472
Neutrophils % 0.0 99.0 69.9 19.59 Blood 79387
Oxygen Saturation % 0.2 901.0 80.0 32.0 Blood 4838
PT sec 8.2 150.0 12.0 3.5 Blood 44656
PTT sec 16.6 150.0 30.6 6.99 Blood 44497
Phosphate mg/dL 0.2 23.8 3.4 1.1 Blood 29888
Platelet Count K/uL 5.0 1898.0 227.0 105.0 Blood 82260
Potassium mEq/L 1.3 10.7 4.2 0.80 Blood 79857
RDW % 10.6 33.8 13.8 2.19 Blood 82266
Red Blood Cells m/uL 0.65 8.06 4.25 0.96 Blood 82266
Sodium mEq/L 88.0 179.0 139.0 5.0 Blood 79865
Troponin T ng/mL 0.0 17.36 0.06 0.13 Blood 41806
Urea Nitrogen mg/dL 1.0 263.0 18.0 13.0 Blood 82522
White Blood Cells K/uL 0.0 632.1 8.1 4.60 Blood 82367
pCO2 mm Hg 8.0 189.0 44.0 17.0 Blood 8008
pH units 5.0 9.0 6.0 1.0 Urine 20968
Table 8: Summary of laboratory values across samples considered in this study, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records for each parameter.
Table 8 provides details on the laboratory values considered in this work, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records. During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence. The specific thresholds for outliers were as follows: absolute basophil count above 20, absolute eosinophil count above 20, absolute lymphocyte count above 100, alanine aminotransferase above 2000, alkaline phosphatase above 2000, aspartate aminotransferase above 2000, creatine kinase above 2000, glucose above 2000, lactate above 2000, and platelet count above 2000. Laboratory values were extracted from the laboratory events table from MIMIC-IV, where for each record we matched all the patient records of the desired tests available between patient record in-time and the 1.5 hours window.
Appendix B Models
Hyperparameter Value
Block of layers 4
S4 model copies 512
S4 state size 8
Optimizer Adam
Learning rate 0.001
Weight decay 0.001
learning rate schedule constant
Batch size 64
Epochs 20
Table 9: S4 model hyperparameters
Table 9 outlines the hyperparameters employed in the S4 model. The architecture consists of four blocks of layers, with each block containing 512 copies of the S4 model. The state size within the S4 model is set to 8. For optimization, the AdamW optimizer is utilized with a learning rate and weight decay both set to 0.001. The learning rate schedule is maintained constant throughout training. A batch size of 64 samples is used for each training iteration, spanning a total of 20 epochs. The training objective is to minimize the binary cross-entropy loss. During training, we apply a model selection strategy on the best performance (AUROC) on the validation set which usually converges earlier than the final number of epochs. As previously described in Appendix A, with the goal of not discarding samples whose any of the deterioration targets happens before the 1.5-hour window or the target event cannot be clearly defined, we introduce a special token in the label space (-999) for which we account during and training and evaluation to be discarded. As of XGboost models, we fit these models with the default hyperparameter values in a binary setting for the later AUROC aggregation across settings, models which in principle do not intrinsically allow the inclusion of samples during training with the special token to be discarded.

### v3

@@
-MDS-ED: Multimodal Decision Support
+a Benchmark Dataset for Diagnoses and Deterioration Prediction in Emergency Medicine
 Abstract
-Background: Benchmarking medical decision support algorithms often struggles due to limited access to datasets, narrow prediction tasks, and restricted input modalities.
-These limitations affect their clinical relevance and performance in high-stakes areas like emergency care, complicating replication, validation, and improvement of benchmarks.
-Methods: We introduce a dataset based on MIMIC-IV, benchmarking protocol, and initial results for evaluating multimodal decision support in the emergency department (ED).
-We use diverse data modalities from the first 1.5 hours of patient arrival, including demographics, biometrics, vital signs, lab values, and electrocardiogram waveforms.
-We analyze 1443 clinical labels across two contexts: predicting diagnoses with ICD-10 codes and forecasting patient deterioration.
-Results: Our multimodal diagnostic model achieves an AUROC score over 0.8 in a statistically significant manner for 357 out of 1428 conditions, including cardiac issues like myocardial infarction and non-cardiac conditions such as renal disease and diabetes.
-The deterioration model scores above 0.8 in a statistically significant manner for 13 out of 15 targets, including critical events like cardiac arrest and mechanical ventilation, ICU admission as well as short- and long-term mortality.
-Incorporating raw waveform data significantly improves model performance, which represents one of the first robust demonstrations of this effect.
-Conclusions: This study highlights the uniqueness of our dataset, which encompasses a wide range of clinical tasks and utilizes a comprehensive set of features collected early during the emergency after arriving at the ED.
-The strong performance, as evidenced by high AUROC scores across diagnostic and deterioration targets, underscores the potential of our approach to revolutionize decision-making in acute and emergency medicine.
+Background: A clinically meaningful comparative assessment of medical decision support in emergency care is challenging due to a lack of appropriate datasets with multimodal input modalities and comprehensive prediction task.
+This hampers measurable progress in the field.
+Results: We introduce a dataset based on MIMIC-IV, a benchmarking protocol, and initial results for evaluating multimodal decision support in the emergency department (ED).
+We use diverse data modalities from the first 1.5 hours after patient arrival, including demographics, biometrics, vital signs, lab values, and electrocardiogram waveforms.
+We analyze 1443 clinical labels across two contexts: predicting diagnoses and patient deterioration.
+Our diagnostic model achieves an AUROC score over 0.8 in a statistically significant manner for 609 out of 1428 conditions, including cardiac conditions like myocardial infarction and non-cardiac conditions such as renal disease and diabetes.
+The deterioration model scores above 0.8 in a statistically significant manner for 14 out of 15 targets, including critical events like cardiac arrest, mechanical ventilation, intensive care unit admission, as well as short- and long-term mortality.
+Furthermore, we provide one of the first robust demonstrations of the significant impact of raw waveform input data on model performance.
+Conclusions: This study highlights the proposed dataset as a unique resource to foster progress towards measurable progress in the domain of algorithmic decision support in emergency care.
+The presented multimodal baseline models showcase the potential of diagnostic decision support in the field and provide strong incentives for including raw waveform data.
+keywords:
+Medical Decision Support; Multimodal Data; Emergency Department; Diagnostic Prediction; Patient Deterioration.
+journal: general
+{keypoints*}
+  • Introduction of a novel dataset based on MIMIC-IV for benchmarking multimodal decision support in the emergency department.
+  • Evaluation of 1443 clinical labels for diagnostic and patient deterioration predictions, achieving high AUROC scores.
+  • Demonstration of significant performance improvements by incorporating raw ECG waveform data.
+  • Highlighting the dataset’s comprehensive feature set collected within the first 1.5 hours after patient arrival.
+  • Potential impact on improving decision-making processes in acute and emergency medicine.
 1 Introduction
@@
 The implementation of AI models is particularly relevant in acute and emergency medicine, where clinicians must address critical conditions within short time frames to make optimal clinical decisions.
-AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8].
+AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8, 9].
 There has been exponential growth in the number of scientific publications on new AI models.
@@
 Benchmarking is crucial for evaluating the performance and reliability of novel AI-models, ensuring their applicability and robustness in real-world clinical settings.
-In this study, we investigate multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
-2 Methods
-2.1 Clinical workflow and dataset creation
-Figure 1: Pipeline outlines the MDS-ED workflow, which involves feature collection encompassing patient demographics, biometrics (such as height, weight, and BMI), vital parameters and trends, laboratory values and trends, and ECG waveform data to address two clinically relevant prediction scenarios: predicting patient discharge diagnoses out of 1428 cardiac and non-cardiac ICD10-CM codes and predicting patient deterioration according to 15 clinical deterioration measures.
-Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay.
-MDS-ED was created by linking ECG waveforms from the MIMIC-IV-ECG [18] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [19].
-Prediction tasks and targets
-In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration.
-For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [11], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG).
-To this end, we frame the task as a multilabel classification task in terms of ICD-10 CM codes.
-As in [11], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits and propagate codes to parent codes up to the third digit.
-In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions.
-For patient deterioration, we build on a consensus definition [20] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA).
-Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days.
-With these two main groups, we aim to cover a wide range spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels, see Table 5 in the supplementary material for a detailed target definition discussion.
-Features
-While aiming to include mainly data at triage, we also include irregularly sampled vital parameters and lab values captured during the 1.5h data collection window.
-To capture trends, we compute simple sample-wise statistical aggregation functions per vital parameter and laboratory values, namely mean, median, minimum, maximum, standard deviation, first and last values, rate of change of them between first and last value, as well as the slope of a linear model fitted on the minutes difference between value collection and arrival as independent variable and the actual values as dependent variables.
-We follow [21] to select an appropriate set of laboratory values that avoids a bias towards particular prediction targets.
-We obtain a total of 470 features across all feature modalities in a tabular format plus a 12-lead 10-second ECG waveform per sample.
-MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,095 samples.
-We leverage the stratified splits provided along with the dataset accompanying [11], which includes stratification based on gender, age bins, and discharge diagnoses.
-We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1, see Appendix A in the supplementary material for further preprocessing and definitions details as well as descriptive statistics on demographics, biometrics, vital parameters and laboratory values in Tables 6-8.
-2.2 Related work
-Group Detail MIMIC-ED-Assist VitalML MEME EHRSHOT MC-BEC MDS-ED
-Source [12] [13] [17] [14] [16] This work
-Population ED ED Longitudinal Longitudinal ED ED
-Size Patients 25714 N/A 947028 6739 63389 71098
-Visits 32356 19847 N/A 921499 102731 121195
-Features Demographics
-Biometrics
-Vital parameters ✓($\mathbb{T}$ ) ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$)
-Lab. values ✓($\mathbb{T}$) ✓($\mathbb{T}$)
-Waveforms $(\mathbb{E}$)
-Chief complaint
-Medications
-Tasks Labels 3 3 4 15 7 1443
-Availability Open source
-Table 1: Direct comparison with related works in terms of dataset size, features, availability, and number of target labels.
-We use diverse symbology to express the contribution where ✓ = available, ✗ = unavailable, $\mathbb{E}$ = available in the form of embeddings, and $\mathbb{T}$ = available in the form of trends or at least two sampled values.
-Table 1 contains a direct comparison with related works in terms of population, dataset size, features, availability, and number of target labels.
-[12] proposed the MIMIC-ED-Assist benchmark to reduce ED length of stay by flagging high-risk patients using triage features.
-[13] proposed VitalML to predict deterioration in the next 90 minutes from the first 15 minutes of monitoring.
-[17] proposed a multiple embedding model for EHR (MEME) to generate pseudo-notes from tabular data and predict ED disposition, discharge location, ICU admission, and inpatient mortality utilizing a natural language processing approach.
-[14] proposed a few shot foundational model EHRSHOT approach to predict operational outcomes, developed diagnoses, as well as anticipate lab test results and image findings.
-[16] proposed MC-BEC to predict patient deterioration, disposition, and revisit from features collected including diagnoses from the first 15 minutes after the patient is assigned to a room.
-Overall, our proposed MDS-ED main contributions in terms of the dataset are: 1) Comprehensive size: MDS-ED is situated in the first position regarding a large number of patients and second in the number of visits in the open-source domain, despite restricting towards the investigated setting of the first 1.5 hours of ED arrival.
-2) Diverse input features: MDS-ED is situated in the first position in terms of feature modalities in the open-source domain as it incorporates a comprehensive set of features, including demographics, biometrics, vital parameter trends, laboratory values trends, and ECG waveforms.
-This is more extensive than most compared datasets.
-Intentionally, we decided to exclude chief complaints as their unstructured nature presents a more challenging step towards external validation with the addition of more variance, see [17].
-Similarly, we exclude previous patient medications as they involve irregularities in administration times, and dosages, as well as potential bias towards task prediction, e.g., certain medications are given for specific gravity and stages of diagnoses.
-3) Comprehensive prediction targets: MDS-ED proposes 1,443 target labels, significantly more than other datasets, which typically cover fewer and too narrow in scope tasks.
-4) Accessibility: Similar to [12, 13, 14], MDS-ED is open source, encouraging further research and collaboration.
-2.3 Model architectures
-To propose an initial benchmark, we experiment with diverse settings based on data modalities input such as (1) tabular only, (2) ECG waveforms only, (3) ECG features + tabular, and (4) ECG waveforms + tabular.
-For waveforms (2) and (4), inspired by recent successful applications in the field of physiological time series [11, 22, 23] we employed an S4 classifier based on structured state space models with four layers [24], specifically, for the multimodal approach in (4) we extended the S4 classifier by fusing it with the tabular features using a tensor fusion approach based on a differentiable outer product proposed in the context of multimodal sentiment analysis [25].
-For (1) and (3) we fit gradient-boosting decision trees (XGBoost).
-Additional details on model architectures and model hyperparameters can be found in Appendix B Table 9.
-2.4 Training and evaluation
-For the S4 models, we used AdamW as optimizer, with both learning rate and weight decay set to 0.001, and maintained a constant learning rate schedule.
-The training was conducted with a batch size of 64 samples over 20 epochs which usually converged earlier and model selection on the validation set.
-The training objective was to minimize binary cross-entropy loss.
-For all the models, we evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC).
-To assess statistical uncertainty resulting from the finite size and specific composition of the test set, we employ empirical bootstrap on the test set with $n=1000$ iterations.
-We report 95% confidence intervals for both macro AUROC and individual label AUROCs.
-Given [11] evaluation protocol, we use only the first record per visit per patient to avoid bias.
-Additional details of training and evaluation procedures can be found in Appendix B.
-3 Results
+In this study, we investigate multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities collected within the first 1.5 hours of patient arrival at the ED department such as demographics, biometrics, vital parameters trends, laboratory trends, and electrocardiogram (ECG) waveforms, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
+2 Analyses
 2.1 Benchmarking predictive performance
+Table 1: Predictive performance (macro AUROC with 95% bootstrap confidence intervals) across all scenarios for unimodal models as well as multimodal prediction models.
+Where the clinical routine data encompasses a comprehensive set of 470 variables from diverse feature modalities such as demographics, biometrics, vital parameters trends, and laboratory values trends.
+All the scenarios can also be split into model architecture, either tree-based or deep-learning models.
 Diagnoses Deterioration
-Tabular only 0.7574 (0.7561, 0.7610) 0.8519 (0.8502, 0.8535)
-ECG waveforms only 0.7631 (0.7590, 0.7673) 0.8279 (0.8047, 0.8511)
-ECG features + Tabular 0.7684 (0.7673, 0.7721) 0.8693 (0.8688, 0.8715)
-ECG waveforms + Tabular 0.7873 (0.7832, 0.7913) 0.8815 (0.8625, 0.8977)
-Table 2: Predictive performance (macro AUROC with 95% bootstrap confidence intervals) across all tasks for unimodal models as well as multimodal prediction models.
+Unimodal models
+Clinical routine data (tree-based) 0.7473 (0.7459, 0.7484) 0.8365 (0.8179, 0.8546)
+ECG features only (tree-based) 0.6366 (0.6352, 0.6374) 0.7385 (0.7114, 0.7588)
+ECG waveforms only (deep-learning) 0.7678 (0.7636, 0.7718) 0.8565 (0.8410 , 0.8707)
+Multimodal models
+ECG features + clinical routine data (tree-based) 0.7685 (0.7676, 0.7705) 0.8656 (0.8498, 0.8830)
+ECG waveforms + clinical routine data (deep-learning) 0.8256 (0.8222,0.8288) 0.9115 (0.8991,0.9222)
 Table 1 presents the predictive performance of both unimodal and multimodal models across the diagnoses and deterioration settings.
-For diagnosis prediction, the multimodal model that integrates ECG waveforms and tabular data achieved an AUROC of 0.7873 (0.7832, 0.7913), which significantly outperformed the unimodal models as well as the ECG features and tabular.
-Interestingly, the unimodal ECG waveforms model outperformed the unimodal tabular model with 0.7631 (0.7590, 0.7673) against 0.7574 (0.7561, 0.7610).
-In the deterioration prediction task, the multimodal model achieved an AUROC of 0.8815 (0.8625, 0.8977), which is notably higher than the unimodal models as well as the ECG features and tabular.
-Interestingly, contrary to the diagnoses setting, the tabular-only model outperformed the ECG waveform-only model with 0.8519 (0.8502, 0.8535) against 0.8279 (0.8047, 0.8511).
-In the following discussion, we restrict ourselves to the discussion of the most comprehensive multimodal model based on tabular features and ECG waveforms.
+For diagnoses prediction, the multimodal model that integrates ECG waveforms and clinical routine data achieved a macro average across all areas under the respective receiver operating curves (AUROC) (macro-AUROC) of 0.8256 (0.8222,0.8288), which is notably higher than the rest of the models.
+Similarly, in the deterioration prediction task, the multimodal model that integrates ECG waveforms and clinical routine data achieved an AUROC of 0.9115 (0.8991,0.9222), which also overcomes the rest of the models.
+In the following discussion, we restrict ourselves to the discussion of the most comprehensive multimodal model based on ECG waveforms and clinical routine data.
 2.2 Task-dependent predictive performance
+Table 2: Best-performing individual statements organized according to selected ICD chapters underscoring the breadth of accurately predictable statements.
+The table shows the four best-performing individual statements per ICD chapter (8 for chapter IX (Circulatory system diseases), 4 for the rest), where we show only AUROC scores above 0.85 and where also the lower bound of the 95% bootstrap confidence interval exceeds 0.80.
+To showcase the breadth of reliably predictable statements, we list only the best-performing statement per 3-digit ICD code.
 Code: AUROC [instances]. Description Code: AUROC [instances]. Description
-IX: Circulatory System. 0.8467 (93/181 >0.80 lower bound AUROC)
-I2119: 0.9975 [395] ST elevation (STEMI) myocardial infarction involving other coronary artery of inferior wall I132: 0.9803 [871] Hypertensive heart and chronic kidney disease with heart failure and with stage 5 chronic kidney disease, or end stage renal disease
-I081: 0.9793 [397] Rheumatic disorders of both mitral and tricuspid valves I447: 0.9691 [1088] Left bundle-branch block, unspecified
-I255: 0.9658 [946] Ischemic cardiomyopathy I420: 0.9616 [315] Dilated cardiomyopathy
-I481: 0.9549 [265] Persistent atrial fibrillation I120: 0.9525 [3350] Hypertensive chronic kidney disease with stage 5 chronic kidney disease or end-stage renal disease
-I30: 0.9514 [316] Acute pericarditis I5023: 0.9494 [3216] Acute on chronic systolic (congestive) heart failure
-III: Blood and immune mechanism: 0.8011 (14/45 >0.80 lower bound AUROC)
-D6181: 0.9333 [1503] Pancytopenia D631: 0.9268 [2960] Anemia in chronic kidney disease
-D6832: 0.9194 [271] Hemorrhagic disorder due to extrinsic circulating anticoagulants D69: 0.8712 [5599] Purpura and other hemorrhagic conditions
-IV: Endocrine, nutritional and metabolic. 0.801 (38/107 >0.80 lower bound AUROC)
-E1143: 0.9529 [249] Type 2 diabetes mellitus with diabetic autonomic (poly)neuropathy E1022: 0.9393 [309] Type 1 diabetes mellitus with diabetic chronic kidney disease
-E43: 0.9216 [1417] Unspecified severe protein-calorie malnutrition E874: 0.9187 [632] Mixed disorder of acid-base balance
-XIV: Genitourinary system 0.7978 (12/39 >0.80 lower bound AUROC)
-N186: 0.9594 [4362] End stage renal disease N08: 0.9334 [798] Glomerular disorders in diseases classified elsewhere
-N170: 0.906 [1773] Acute kidney failure with tubular necrosis N2581: 0.8816 [472] Secondary hyperparathyroidism of renal origin
-XIX: Injury, poisoning: 0.797 (20/84 >0.80 lower bound AUROC)
-S0180: 0.9143 [261] Unspecified open wound of other part of head S063: 0.9125 [249] Focal traumatic brain injury
-T8612: 0.9075 [559] Kidney transplant failure T4551: 0.9021 [572] Poisoning by, adverse effect of and underdosing of anticoagulants
-II: Neoplasms. 0.792 (6/53 >0.80 lower bound AUROC)
-C228: 0.9326 [236] Malignant neoplasm of liver, primary, unspecified as to type C786: 0.9198 [490] Secondary malignant neoplasm of retroperitoneum and peritoneum
-C83: 0.8825 [370] Non-follicular lymphoma C7952: 0.8794 [629] Secondary malignant neoplasm of bone marrow
-I: Infectious and parasitic diseases 0.7908 (7/51 >0.80 lower bound AUROC)
-A4189 0.9688 [267] Other specified sepsis B9681 0.8749 [245] Helicobacter pylori [H. pylori] as the cause of diseases
-X: Respiratory system. 0.7886 (23/71 >0.80 lower bound AUROC)
-J9691 0.9530 [572] Respiratory failure, unspecified with hypoxia J152 0.9438 [289] Pneumonia due to staphylococcus
-J841 0.9136 [353] Other interstitial pulmonary diseases with fibrosis J948 0.9048 [369] Other specified pleural conditions
-XXI: Health status and services 0.7884 (40/163 >0.80 lower bound AUROC)
-Z992 0.9681 [3224] Dependence on renal dialysis Z902 0.9633 [265] Acquired absence of lung [part of]
-Z681 0.9535 [853] Body mass index (BMI) 19.9 or less, adult Z4502 0.9533 [345] Encounter for adjustment and management of automatic implantable cardiac defibrillator
-XI: Digestive system 0.7825 (29/119 >0.80 lower bound AUROC)
-K7031: 0.9817 [485] Alcoholic cirrhosis of liver with ascites K7290: 0.9605 [548] Hepatic failure, unspecified without coma
-K758: 0.9299 [457] Other specified inflammatory liver diseases K762: 0.9272 [350] Central hemorrhagic necrosis of liver
-XII: Skin and subcutaneous tissue 0.7804 (5/25 >0.80 lower bound AUROC)
-L9750: 0.9303 [297] Non-pressure chronic ulcer of foot L409: 0.8694 [239] Psoriasis
-V: Mental and behavioral disorders 0.7719 (11/71 >0.80 lower bound AUROC)
-F500: 0.9856 [262] Anorexia nervosa F1012: 0.9243 [625] Alcohol abuse with intoxication
-F4310: 0.9050 [748] Post-traumatic stress disorder, unspecified F1210: 0.8753 [396] Cannabis abuse, uncomplicated
-XVIII: Symptoms, signs, and abnormal findings 0.7477 (19 /167 >0.80 lower bound AUROC)
-R6521: 0.9561 [2545] Severe sepsis with septic shock R5081: 0.9532 [406] Fever presenting with conditions classified elsewhere
-R570: 0.9507 [981] Cardiogenic shock R297: 0.9273 [290] National Institutes of Health Stroke Scale (NIHSS) score
-XIII: Musculoskeletal system 0.7282 (8/66 >0.80 lower bound AUROC)
-M6282: 0.9224 [598] Rhabdomyolysis M7966: 0.8944 [231] Pain in lower leg
-M8580: 0.8931 [540] Other specified disorders of bone density and structure M86 0.8838 [489] Osteomyelitis
-Table 3: Best-performing individual statements organized according to selected ICD chapters underscoring the breadth of accurately predictable statements.
-The table shows the four best-performing individual statements per ICD chapter (10 for chapter IX (Circulatory system diseases), 4 for the rest), where we show only AUROC scores above 0.85 and where also the lower bound of the 95% bootstrap confidence interval exceeds 0.80.
-To showcase the breadth of reliably predictable statements, we list only the best-performing statement per 3-digit ICD code.
+IX: Circulatory system. 0.8761 (126/181 >0.80 lower bound)
+I071: 0.9461 [287] Rheumatic tricuspid insufficiency I080: 0.8522 [770] Rheumatic mitral and aortic valve disorders
+I110: 0.9044 [3185] Hypertensive heart disease with heart failure I120: 0.9756 [3349] Hypert. CKD (stage 5 or with end-stage renal disease)
+I130: 0.9426 [2994] Hypertensive heart and kidney disease with heart failure and CKD I160: 0.9524 [427] Hypertensive urgency
+I200: 0.8934 [810] Unstable angina I2109: 0.9734 [401] STEMI anterior wall coronary artery
+II: Neoplasms. 0.8572 (34/53 >0.80)
+C228: 0.9628 [235] Malignant neoplasm of liver, primary C259: 0.8991 [275] Malignant neoplasm of pancreas
+C341: 0.8739 [463] Malignant neoplasm of upper lobe or lung C5091: 0.8544 [425] Malignant neoplasm of breast, female
+XIX: Injury, poisoning, and certain other consequences of external causes. 0.8479 (37/86 >0.80)
+S0100: 0.9161 [258] Unspecified open wound of scalp S065: 0.8767 [783] Traumatic subdural hemorrhage
+S12: 0.9356 [392] Fracture of cervical vertebra and other parts of neck S2249: 0.9198 [409] Multiple fractures of ribs
+I: Certain infectious and parasitic diseases. 0.8346 (19/52 >0.80)
+A0472: 0.9374 [235] Enterocolitis due to Clostridium difficile A084: 0.8709 [352] Viral intestinal infection,
+A40: 0.9347 [339] Streptococcal sepsis A4151: 0.9522 [626] Sepsis due to Escherichia coli [E. coli]
+IV: Endocrine, nutritional, and metabolic diseases. 0.8336 (52/107 >0.80)
+E1022: 0.9661 [309] Type 1 diabetes mellitus with diabetic CKD E1121: 0.9466 [707] Type 2 diabetes mellitus with diabetic nephropathy
+E1342: 0.9271 [1809] Diabetes mellitus with diabetic polyneuropathy E43: 0.9329 [1417] Severe protein-calorie malnutrition
+III: Diseases of the blood and blood-forming organs. 0.8335 (20/45 >0.80)
+D500: 0.8557 [897] Cronic Iron deficiency anemia D6181: 0.8761 [1503] Pancytopenia
+D62: 0.8843 [4514] Acute posthemorrhagic anemia D630: 0.8920 [589] Anemia in neoplastic disease
+XI: Diseases of the digestive system. 0.827 (40/119 >0.80)
+K264: 0.9189 [255] Chronic duodenal ulcer with hemorrhage K3181: 0.9384 [419] Angiodysplasia of stomach and duodenum
+K521: 0.9290 [257] Toxic gastroenteritis and colitis K566: 0.8958 [598] Other and unspecified intestinal obstruction
+XIV: Diseases of the genitourinary system. 0.8253 (18/39 >0.80)
+N08: 0.9393 [798] Glomerular disorders N12: 0.8941 [325] Tubulo-interstitial nephritis, not acute or chronic
+N170: 0.9092 [1773] Acute kidney failure with tubular necrosis N183: 0.8625 [4734] Chronic kidney disease, stage 3
+X: Diseases of the respiratory system. 0.8226 (33/71 >0.80)
+J1521: 0.9202 [275] Pneumonia due to staphylococcus aureus J39: 0.9144 [276] Other diseases of upper respiratory tract
+J440: 0.8936 [440] Chronic obstructive pulmonary disease with acute lower respiratory infection J690: 0.8968 [2231] Pneumonitis due to inhalation of food and vomit
+V: Mental, behavioral and neurodevelopmental disorders. 0.8111 (32/71 >0.80)
+F0280: 0.8556 [1540] Dementia F0390: 0.8565 [3690] Unspecified dementia
+F1012: 0.9591 [625] Alcohol abuse with intoxication F1110: 0.8735 [666] Opioid abuse, uncomplicated
+XII: Diseases of the skin and subcutaneous tissue. 0.8087 (10/25 >0.80)
+L0311: 0.8646 [573] Cellulitis of other parts of limb L408: 0.8933 [292] Other psoriasis
+L8915: 0.8920 [1209] Pressure ulcer of sacral region L974: 0.9093 [336] Non-pressure chronic ulcer of heel and midfoot
+VI: Diseases of the nervous system. 0.7955 (19/67 >0.80)
+G309: 0.8508 [1192] Alzheimer’s disease G4030: 0.9374 [244] Idiopathic epilepsy and epileptic syndromes
+G609: 0.8772 [816] Hereditary and idiopathic neuropathy G8190: 0.9251 [265] Hemiplegia
+XVIII: Symptoms, signs and abnormal clinical and laboratory findings 0.7813 (35/171 >0.80)
+R000: 0.8507 [2714] Tachycardia, unspecified R042: 0.8745 [316] Hemoptysis
+R0600: 0.8502 [1999] Dyspnea R1032: 0.8956 [453] Left lower quadrant pain
+XIII: Diseases of the musculoskeletal system and connective tissue. 0.7405 (14/68 >0.80)
+M100: 0.9106 [246] Idiopathic gout M129: 0.8933 [495] Arthropathy
+M1A9: 0.9000 [253] Chronic gout, unspecified M810: 0.8439 [4398] Age-related osteoporosis without fracture
 Table 2 demonstrates the diagnoses that can be predicted most accurately by the model, which we organized based on ICD chapters.
-The algorithm’s predictive performance ranges from an AUROC of 0.7282 for the Musculoskeletal System (XIII) chapter to 0.8467 for the Circulatory System (IX) chapter.
-In total, the model predicted 325 out of 1,242 individual ICD diagnoses with high accuracy, which we define in this work as conditions where the lower bound of the 95% bootstrap confidence interval for AUROC exceeding 0.80.
+The algorithm’s predictive performance ranges from an AUROC of 0.7405 for the musculoskeletal system and connective tissue (XIII) chapter to 0.8761 for the circulatory system (IX) chapter.
+In total, the model predicted 609 out of 1,428 individual ICD diagnoses with high accuracy, which we define in this work as conditions where the lower bound of the 95% bootstrap confidence interval for AUROC exceeds 0.80.
+Table 3: Detailed breakdown of model performance in the deterioration category.
 Label: AUROC [instances] Label: AUROC [instances]
-Clinical deterioration. 0.8620 (4/6 >0.80 lower bound AUROC)
-Severe hypoxemia: 0.5839 [555] ECMO: 0.8705 [166]
-Vasopressors: 0.9132 [1172] Inotropes: 0.9079 [376]
-Mechanical ventilation: 0.9394 [4230] IHCA: 0.9576 [623]
-ICU admission. 0.885 (2/2 >0.80 lower bound AUROC)
-24-hours: 0.8923 [15688] Overall: 0.8777 [18872]
-Mortality. 0.8970 (7/7 >0.80 lower bound AUROC)
-In-hospital: 0.8954 [99] 24-hours: 0.9512 [817]
-7-days: 0.9149 [2275] 28-days: 0.9009 [5153]
-90-days: 0.8850 [9300] 180-days: 0.8692 [12844]
-365-days: 0.8629 [17761]
-Table 4: Detailed breakdown of model performance in the deterioration category.
+Clinical deterioration. 0.9070 (5/6 > 0.80) lower bound
+Severe hypoxemia: 0.6980 [555] ECMO: 0.9355 [166]
+Vasopressors: 0.9239 [1171] Inotropes: 0.9400 [375]
+Mechanical ventilation: 0.9590 [4223] IHCA: 0.9859 [623]
+ICU admission. 0.9063 (2/2 > 0.80)
+ICU 24 hours: 0.9147 [15677] ICU overall: 0.8979 [18860]
+Mortality. 0.9168 (7/7 > 0.80)
+In-hospital: 0.9423 [99] 24 hours: 0.9600 [816]
+7 days: 0.9429 [2273] 28 days: 0.9115 [5150]
+90 days: 0.8952 [9297] 180 days: 0.8894 [12839]
+365 days: 0.8768 [17756]
 Table 3 showcases the deterioration model’s predictive performance across clinical deterioration, ICU admission, as well as both short and long-term mortality with many AUROC scores significantly exceeding the 0.80 threshold.
-Overall, for clinical deterioration, the model achieves an AUROC of 0.8692, whereas for ICU admissions, the model reports an overall AUROC of 0.8842, and for mortality predictions also exhibits high performance, with an overall AUROC of 0.9030.
+Overall, for clinical deterioration, the model achieves an AUROC of 0.9070, whereas, for ICU admissions, the model reports an overall AUROC of 0.9063, and for mortality predictions also exhibits high performance, with an overall AUROC of 0.9168.
 3 Discussion
 3.1 Impact of data modalities
-From Section 3.1 we can draw several conclusions: Firstly, the results demonstrate that multimodal models, which integrate diverse data types, offer superior performance in both diagnostic and deterioration tasks (row 1/2 vs.
-row 3/4).
-Secondly, in the diagnoses task, the use of ECG raw waveforms instead of ECG features improves the performance in a statistically significant manner (row 3 vs.
-row 4), whereas for the deterioration task, we performed a direct comparison via bootstrapping the score difference for statistical significance, and we did not find a statistically significant difference, which is in line with [13].
-To the best of our knowledge, this is the first statistically robust demonstration of the added value of raw waveform input against features for clinically relevant prediction tasks such as diagnoses prediction.
-Thirdly, the model building on ECG waveforms as only input outperforms the tabular-only model in the diagnostic setting, but not in the deterioration setting (row 1 vs.
-row 2).
-We hypothesize that this is due to the inclusion of tabular trends over time which aligns with the task definition of deterioration.
-We believe that the inclusion of multiple raw ECGs over time instead of just a single snapshot would allow us to capture more meaningful deterioration trends also from raw waveform data.
+From the results presented in the previous section, we can draw several conclusions: Firstly, the results demonstrate that multimodal models, which integrate diverse data types, offer superior performance in both diagnostic and deterioration tasks (row 4&5 vs.
+the rest).
+Secondly, in the diagnoses task as well as in the deterioration task, the use of ECG raw waveforms instead of ECG features improves the performance in a statistically significant manner (row 4 vs.
+row 5), finding which is not in line with [14].
+To the best of our knowledge, this is the first statistically robust demonstration of the added value of raw ECG waveform input against ECG features for clinically relevant prediction tasks such as diagnoses and deterioration prediction.
+Thirdly, for the unimodal models, in the deterioration task, the clinical routine data model outperforms ECG features only and ECG waveforms only, however, in the diagnoses task, the ECG waveforms only outperforms the other 2 settings, we hypothesize that for the deterioration task, the clinical routine data apart of including a rich set of clinical features (demographics, biometrics, vital parameters trends, and laboratory values trends) against only an single ECG either in features or waveform, it also includes trends over time which aligns with the task definition of deterioration.
+Despite this, we believe that a single ECG snapshot can achieve high performances for both tasks, but also we believe that the inclusion of multiple ECGs over time instead of just a single snapshot would allow us to capture more meaningful deterioration and potentially diagnoses trends.
 3.2 Clinical significance
-From Section 3.2, we conclude that our diagnostic models exhibit high predictive capabilities for both acute and chronic conditions.
+From a clinical point of view, we conclude that our diagnostic models exhibit high predictive capabilities for acute, chronic, as well as management conditions.
 Acute conditions like ST elevation myocardial infarction, acute pericarditis, persistent atrial fibrillation, and acute kidney failure require rapid diagnosis and immediate intervention to prevent severe outcomes.
@@
 The model’s ability to effectively predict these chronic conditions underscores its utility in supporting sustained patient monitoring and management, ultimately improving patient outcomes through consistent and reliable diagnostic support.
+Finally, detecting management conditions like thrombosis of prosthetic devices enhances ongoing care and ensures timely intervention.
 As demonstrated in [12], even the model based only on ECG waveforms can reliably predict cardiac but most notably non-cardiac conditions.
@@
 Additionally, its robust performance in predicting ICU admissions allows for better resource allocation, ensuring that patients who need intensive care receive it promptly while preventing unnecessary admissions.
-The model accurately predicts 24-hour in-hospital mortality and has high accuracy in predicting 365-day mortality, supporting clinical decision-making in prioritizing high-risk patients and discussing prognosis with families.
+The model accurately predicts 24-hour in-hospital mortality and accurately predicts 365-day mortality, supporting clinical decision-making in prioritizing high-risk patients and discussing prognosis with families.
 Long-term mortality predictions aid in advanced care planning, such as deciding against ICU admission for patients with a short expected lifespan after discharge, and inform post-discharge planning and follow-up care to prevent readmissions and improve long-term health outcomes.
@@
 associational attributions along the lines of [21] is beyond the scope of this submission.
-Finally, although the open-source dataset fosters innovation, the implementation of a large-scale single end-to-end model with a wide range of target variables such as the diagnoses case might result in challenges during deployment in resource-constrained hardware, see Appendix B in the supplementary material for a more extensive discussion.
+Finally, although the open-source dataset fosters innovation, the implementation of a large-scale single end-to-end model with a wide range of target variables such as the diagnoses case might result in challenges during deployment in resource-constrained hardware, see 9 in the supplementary material for a more extensive discussion.
 3.4 Future research directions
@@
 Going beyond MIMIC-IV, models developed on the proposed dataset could foreshadow the development of more comprehensive models on more specialized datasets such as [22], for example including molecular data as input and/or prediction targets, to address more fine-grained diagnostic or deterioration prediction targets.
-4.5 Social impact
+4 Potential implications
 Clinically, these models enhance decision-making by providing timely and accurate predictions, reducing diagnostic errors, and improving patient care, especially in high-stakes emergency settings.
@@
 However, questions of algorithmic fairness remain to be investigated in detail in future work.
-4.6 Authors contribution statement
+5 Methods
+5.1 Clinical workflow and dataset creation
+Figure 1: Pipeline outlines the MDS-ED workflow, which involves feature collection encompassing patient demographics, biometrics (such as height, weight, and BMI), vital parameters and trends, laboratory values and trends, and ECG waveform data to address two clinically relevant prediction scenarios: predicting patient discharge diagnoses out of 1428 cardiac and non-cardiac ICD10-CM codes and predicting patient deterioration according to 15 clinical deterioration measures.
+Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay.
+MDS-ED was created by linking ECG waveforms from the medical information mart for intensive care (MIMIC) database, specifically, MIMIC-IV-ECG [23] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [24].
+Prediction tasks and targets
+In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration.
+For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [12], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG).
+To this end, we frame the task as a multilabel classification task in terms of international classification of diseases clinical modification 10 (ICD-10 CM) codes.
+As in [12], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits, and propagate codes to parent codes up to the third digit.
+In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions.
+For patient deterioration, we build on a consensus definition [25] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA).
+Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days.
+With these two main groups, we aim to cover a wide range spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels, see Table 5 in the supplementary material for a detailed target definition discussion.
+Features
+While aiming to include mainly data at triage, we also include irregularly sampled vital parameters and lab values captured during the 1.5h data collection window.
+To capture trends, we compute simple sample-wise statistical aggregation functions per vital parameter and laboratory values, namely mean, median, minimum, maximum, standard deviation, first and last values, rate of change of them between first and last value, as well as the slope of a linear model fitted on the minutes difference between value collection and arrival as independent variable and the actual values as dependent variables.
+We follow [26] to select an appropriate set of laboratory values that avoids a bias towards particular prediction targets.
+We obtain a total of 470 features of clinical routine data across all feature modalities in a tabular format plus a 12-lead 10-second ECG waveform per sample.
+Similarly, the ECG feature set is composed of RR interval, P-onset, QRS-onset, QRS-end, and T-end in milliseconds (ms), whereas P-axis, QRS-axis, and T-axis in degrees.
+MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,057 samples.
+We leverage the stratified splits provided along with the dataset accompanying [12], which includes stratification based on gender, age bins, and discharge diagnoses.
+We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1.
+For deep learning models with clinical routine data, we apply a median imputation learned from train set to validation and test for each column and we apply a masking strategy in which we add an additional binary column to indicate which values were imputed to the model.
+See the supplementary material for further preprocessing and definitions details as well as descriptive statistics on demographics, biometrics, vital parameters and laboratory values in Tables 6-8.
+5.2 Related work
+Table 4: Direct comparison with related works in terms of dataset size, features, availability, and number of target labels.
+Group Detail MIMIC-ED-Assist VitalML MEME EHRSHOT MC-BEC MDS-ED
+Source [13] [14] [18] [15] [17] This work
+Population ED ED Longitudinal Longitudinal ED ED
+Size Patients 25714 N/A 947028 6739 63389 71098
+Visits 32356 19847 N/A 921499 102731 121195
+Features Demographics
+Biometrics
+Vital parameters ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$)
+Lab. values ✓($\mathbb{T}$) ✓($\mathbb{T}$)
+Waveforms ✓($\mathbb{E}$)
+Chief complaint
+Medications
+Tasks Labels 3 3 4 15 7 1443
+Availability Open source
+{tablenotes}
+We use diverse symbology to express the contribution where ✓ = available, ✗ = unavailable, $\mathbb{E}$ = available in the form of embeddings, and $\mathbb{T}$ = available in the form of trends or at least two sampled values.
+Table 4 contains a direct comparison with related works in terms of population, dataset size, features, availability, and number of target labels.
+[13] proposed the MIMIC-ED-Assist benchmark to reduce ED length of stay by flagging high-risk patients using triage features.
+[14] proposed VitalML to predict deterioration in the next 90 minutes from the first 15 minutes of monitoring.
+[18] proposed a multiple embedding model for electronic health records (EHR) (MEME) to generate pseudo-notes from tabular data and predict ED disposition, discharge location, ICU admission, and inpatient mortality utilizing a natural language processing approach.
+[15] proposed a few shot foundational model EHRSHOT approach to predict operational outcomes, developed diagnoses, as well as anticipate lab test results and image findings.
+[17] proposed MC-BEC to predict patient deterioration, disposition, and revisit from features collected including diagnoses from the first 15 minutes after the patient is assigned to a room.
+Overall, our proposed MDS-ED main contributions in terms of the dataset are: 1) Comprehensive size: MDS-ED is situated in the first position regarding a large number of patients and second in the number of visits in the open-source domain, despite restricting towards the investigated setting of the first 1.5 hours of ED arrival.
+2) Diverse input features: MDS-ED is situated in the first position in terms of feature modalities in the open-source domain as it incorporates a comprehensive set of features, including demographics, biometrics, vital parameter trends, laboratory values trends, and ECG waveforms.
+This is more extensive than most compared datasets.
+Intentionally, we decided to exclude chief complaints as their unstructured nature presents a more challenging step towards external validation with the addition of more variance, see [18].
+Similarly, we exclude previous patient medications as they involve irregularities in administration times, and dosages, as well as potential bias towards task prediction, e.g., certain medications are given for specific gravity and stages of diagnoses.
+3) Comprehensive prediction targets: MDS-ED proposes 1,443 target labels, significantly more than other datasets, which typically cover fewer and too narrow in scope tasks.
+4) Accessibility: Similar to [13, 14, 15], MDS-ED is open source, encouraging further research and collaboration.
+5.3 Model architectures
+To propose an initial benchmark, we experiment with diverse settings based on data modalities input such as (1) clinical routine data, (2) ECG features only, (3) ECG waveforms only, (4) ECG features + clinical routine data, and (4) ECG waveforms + clinical routine data.
+For waveforms (3) and (5), inspired by recent successful applications in the field of physiological time series [12, 27, 28] we employed an S4 classifier based on structured state space models with four layers [29] (with state dimension 8 and model dimension 512 as in [28]).
+For the multimodal models in (5), we use a 3-layer MLP as tabular encoder, where the three categorical features (gender, race, acuity score) are first processed through appropriate embedding layers.
+The outputs of the signal encoder and the tabular encoder are then fused through concatenation.
+For (1), (2) and (4) we fit extreme gradient boosting (XGBoost) decision trees.
+Additional details on model architectures and model hyperparameters can be found in 9 Table 9.
+5.4 Training and evaluation
+For the S4 models, we used Adam Weight Decay Optimization (AdamW) as the optimizer, with both learning rate and weight decay set to 0.001, and maintained a constant learning rate schedule.
+The training was conducted with a batch size of 64 samples over 20 epochs which usually converged earlier and model selection on the validation set.
+The training objective was to minimize binary cross-entropy loss.
+For all the models, we evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC).
+To assess statistical uncertainty resulting from the test set’s finite size and specific composition, we employ empirical bootstrap on the test set with $n=1000$ iterations.
+We report 95% confidence intervals for both macro AUROC and individual label AUROCs.
+Given [12] evaluation protocol, we use only the first record per visit per patient to avoid bias.
+Additional details of training and evaluation procedures can be found in 9.
+6 Availability of source code and requirements
+  • Project name: MDS-ED
+  • Project home page: [30].
+  • Project code repository: [31].
+  • Operating system(s): Platform independent
+  • Programming language: Python
+  • Other requirements: Python 3.10.8, PyTorch 1.13.0, and PyTorch Lightening 1.8.0.
+  • License: MIT
+7 Declarations
+7.1 List of abbreviations
+AdamW: Adam Weight Decay Optimization, AI: Artificial Intelligence, AUROC: Area Under the Receiver Operating Characteristic Curve, ECG: Electrocardiogram, ECMO: Extracorporeal Membrane Oxygenation, ED: Emergency Department, EHR: Electronic Health Records, ICD CM: International Classification of Diseases Clinical Modification, ICU: Intensive Care Unit, IHCA: In-Hospital Cardiac Arrest, MDS-ED: Multimodal Decision Support for Emergency Department, MIMIC: Medical Information Mart for Intensive Care, MLP: multi-layer perceptron, XGBoost: Extreme Gradient Boosting.
+7.2 Consent for publication
+Not applicable
+7.3 Competing Interests
+The authors have declared no competing interests.
+7.4 Author’s Contributions
 JMLA and NS conceived and designed the project.
-JMLA conducted the full experimental analyses, with NS supervising them, and HB providing critical revision for clinical intellectual content.
+NS and JMLA implemented the prediction models.
+JMLA conducted the experimental analyses, with NS supervising them, and HB providing critical revision for clinical intellectual content.
 JMLA produced the first draft, NS and HB revised it.
 All authors critically revised the content and approved the final version for publication.
-4.7 Code and data availability
-Code for experiment reproducibility and dataset instructions in a dedicated code repository [30]. For easier usage, the dataset is currently prepared for submission as a Physionet dataset.
-Appendix A Dataset details
-Name Labels Details
-Diagnoses model (1428 labels, based on [11])
-Diagnoses 1428 Discharge diagnoses ICD10-CM codes
-Deterioration model (15 labels, based on [20])
-Mortalities 6 1, 7, 30, 90, 180, 365 days and in-stay
-ICU Admission 2 24 hours and in-stay
-Vasopressors 1 Within 24 hours: epinephrine, norepinephrine, vasopressin, dobutamine, dopamine, or phenylephrine
-Inotropes 1 Within 24 hours: epinephrine, dobutamine, or dopamine
-Mechanical Ventilation 1 Same day or next day ICD9/ICD10 codes: 9670, 9671, 9672, 5A1935Z, 5A1945Z, and 5A1955Z
-In-Hospital Cardiac Arrest (IHCA) 1 Same day or next day ICD9/ICD10 codes: I469, 4275, I462, V1253, I468
-Extracorporeal Membrane Oxygenation (ECMO) 1 Same day or next day ICD9/ICD10 codes: 3961, 3965, 3966, 5A1221Z, 5A1522G, 5A1522H, 5A15223, 5A1522F, 5A15A2F, 5A15A2G, and 5A15A2H
-Severe Hypoxemia 1 Oxygen saturation $\leq$ 85% within 24 hours
-Table 5: Definitions of labels for the diagnoses and deterioration models, including 1428 discharge diagnoses based on ICD10-CM codes in the diagnoses model, and 15 labels covering clinical outcomes such as mortalities, ICU admission, vasopressors, inotropes, mechanical ventilation, in-hospital cardiac arrest (IHCA), extracorporeal membrane oxygenation (ECMO), and severe hypoxemia in the deterioration model.
-Table 5 contains the considered tasks by description, number of labels, and additional details such as target criteria definition.
-Apart from diagnoses, we use the rest of the tasks defined as patient deterioration based on previous work from the literature that defines physiological deterioration [20].
-Although we consider diverse scores such as SOFA significant, calculating these markers requires a more detailed definition of downstream task prediction, which goes beyond the scope of maintaining homogeneity across tasks in this study.
-The curation of the dataset starts from MIMIC-IV-ECG-ICD [11], from which we obtained the ICD10 diagnoses codes for the complete MIMIC-IV-ECG [18] dataset, therefore, based on ECG records with available ED stay_id, we select the ones that happened between the patient in time and the 1.5 hours window.
-Mortalities at finite horizons were computed based on the date of death from [18], and for the computation of mortality per stay we include discharge date times from the admissions table in [19] MIMIC-IV to account for final discharge, all of the mortalities dates were previously validated from the source with a follow-up up to 1 year after the patient’s last recorded discharge.
-ICU admission in 24 hours and per stay were computed based on the ICU stays table from [19].
-Mechanical ventilation and ECMO were computed given the procedures table from MIMIC-IV-ED given the patient admission and procedure date.
-Cardiac arrest was computed from the diagnosis tables from MIMIC-IV-ED and MIMIC-IV, where we first validated if the patient was discharged with the diagnosis code within the ED or hospital otherwise.
-We clarify that although mechanical ventilation, ECMO, and cardiac arrest do not provide an exact time of events, we carefully approximate the times with dates given the same day or the day after.
-Vasopressors and inotropes were computed from the pyxis table from MIMIC-IV-ED given the exact time of the applied medication.
-Finally, severe hypoxemia was computed from the vital signs table from MIMIC-IV-ED.
-In the deterioration setting, to not discard samples where the events happened before the 1.5 hours for a single task, e.g.
-severe hypoxemia within 1.5 hours, we include a special token for the task in the label space which we mask during training and evaluation in order not to negatively affect training and evaluation for other labels due to exclusion of entire patients, see Appendix B for more details.
-Name Values
-Gender (%)
-Female 68289 (52.9)
-Male 60806 (47.1)
-Age (%)
-Median years (SD) 64 (28)
-Quantile 1: 18-49 32634 (25.27)
-Quantile 2: 50-64 34376 (26.62)
-Quantile 3: 65-77 31821 (24.64)
-Quantile 4: 78-101 30264 (23.44)
-Ethnicity (%)
-White 78310 (60.66)
-Black 27680 (21.44)
-Hispanic 9230 (7.14)
-Asian 4700 (3.64)
-Other 9175 (7.10)
-Table 6: Summary demographic characteristics across samples including gender distribution, age distribution (median with standard deviation and quartiles), and ethnicity distribution in the study population, presented as percentages.
-Table 6 provides detailed demographic information on the study population, including gender, age, and ethnicity.
-The gender distribution shows a slightly higher proportion of females (52.9%) compared to males (47.1%).
-The age distribution is divided into four quantiles, with a median age of 64 years and a standard deviation of 28 years.
-Regarding ethnicity, the majority of the population is White (60.66%), followed by Black (21.44%), Hispanic (7.14%), Asian (3.64%), and individuals of other ethnicities (7.10%).
-This diverse demographic distribution ensures a comprehensive evaluation of the model across various patient groups.
-These variables were extracted from the ed stays table from MIMIC-IV-ED.
-Name Unit Minimum Maximum Median IQR Records
-Biometrics
-BMI kg/m2 0 99.2 27.7 8.70 70321
-Weight kg 20.41 367.99 77.11 27.56 73646
-Height cm 64.99 387.09 167.64 15.23 32573
-Vital parameters
-temperature C 13.3 44.72 36.6 0.55 77179
-heartrate bpm 0.0 570.0 84.0 30.0 153249
-resprate bpm 0.0 169.0 18.0 4.0 150525
-o2sat % 0.0 100.0 99.0 3.0 146140
-sbp mmHg 0.0 274.0 130.0 35.0 151346
-dbp mmHg 0.0 486.0 73.0 23.0 151254
-Triage
-Acuity level 1 5 2 1 126071
-Table 7: Summary of statistics across samples for biometric measurements, vital parameters, and triage acuity levels, including minimum, maximum, median, interquartile range (IQR), and number of records for each variable.
-Table 7 provides details on the biometrics, vital parameters, and triage values considered in this work, including the name, unit, minimum value, maximum value, median value, and the number of records.
-Dataset preprocessing includes weight and height conversion from pounds and inches to kilograms and centimeters.
-During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence.
-The specific thresholds for outliers were as follows: temperature below and above 50 and 150 Fahrenheit degrees respectively.
-Heart rate above 700, respiration rate above 300, oxygen saturation below 0 and above 100, diastolic and systolic blood pressures above 500.
-BMI above 100, weight below 20 or above 400, and height below 60 and above 400.
-The temperature unit was converted from Fahrenheit to Celsius, whereas weight from pounds to kilograms, and height from inches to centimeters.
-Biometrics features were extracted from the online medical records table from MIMIC-IV from which we match the closest value of a patient stay within 30 days of admission, otherwise missing.
-Vital parameters and acuity were extracted from the vital signs and triage tables of MIMIC-IV-ED.
-Name Unit Minimum Maximum Median IQR Fluid Records
-Abs. Basophil Count K/uL 0.0 9.2 0.04 0.03 Blood 36837
-Abs. Eosinophil Count K/uL 0.0 15.33 0.09 0.15 Blood 36838
-Abs. Lymphocyte Count K/uL 0.0 88.08 1.47 1.12 Blood 36909
-Alanine Aminotransf. IU/L 1.0 1999.0 22.0 22.0 Blood 29999
-Albumin g/dL 1.0 6.7 4.0 0.79 Blood 28324
-Alkaline Phosphatase IU/L 5.0 1985.0 85.0 51.0 Blood 29923
-Asparate Aminotransf. IU/L 1.0 1996.0 29.0 29.0 Blood 29995
-Bands % 0.0 61.0 0.0 3.0 Blood 6057
-Base Excess mEq/L -413.0 29.0 0.0 6.0 Blood 8011
-Basophils % 0.0 35.0 0.4 0.39 Blood 79386
-Bicarbonate mEq/L 2.0 50.0 24.0 5.0 Blood 79771
-Bilirubin, Direct mg/dL 0.0 45.5 1.0 2.5 Blood 1174
-Bilirubin, Total mg/dL 0.0 55.8 0.5 0.5 Blood 30098
-C-Reactive Protein mg/L 0.1 299.3 17.95 75.14 Blood 1108
-Calcium, Total mg/dL 0.0 79.2 9.2 0.79 Blood 30262
-Carboxyhemoglobin % 0.0 18.0 2.0 2.0 Blood 1078
-Chloride mEq/L 58.0 140.0 102.0 6.0 Blood 79855
-Creatine Kinase (CK) IU/L 8.0 1998.0 115.0 158.0 Blood 9588
-CK, MB Isoenzyme ng/mL 1.0 497.0 3.0 4.0 Blood 9272
-Creatinine mg/dL 0.0 43.0 1.0 0.5 Blood 82850
-Eosinophils % 0.0 75.0 1.3 2.1 Blood 79387
-Fibrinogen, Functional mg/dL 35.0 1348.0 298.5 161.5 Blood 2280
-Free Calcium mmol/L 0.48 3.31 1.08 0.12 Blood 1472
-Glucose mg/dL 3.8 1684.0 113.0 50.0 Blood 85022
-Hematocrit % 6.9 66.1 38.3 7.89 Blood 82490
-Hemoglobin g/dL 0.0 21.8 12.6 3.0 Blood 84890
-INR(PT) nan 0.75 24.0 1.1 0.30 Blood 44656
-Lactate mmol/L 0.3 25.2 1.8 1.20 Blood 32666
-Lymphocytes % 0.0 100.0 19.7 17.1 Blood 79387
-Magnesium mg/dL 0.2 45.0 2.0 0.40 Blood 30472
-Neutrophils % 0.0 99.0 69.9 19.59 Blood 79387
-Oxygen Saturation % 0.2 901.0 80.0 32.0 Blood 4838
-PT sec 8.2 150.0 12.0 3.5 Blood 44656
-PTT sec 16.6 150.0 30.6 6.99 Blood 44497
-Phosphate mg/dL 0.2 23.8 3.4 1.1 Blood 29888
-Platelet Count K/uL 5.0 1898.0 227.0 105.0 Blood 82260
-Potassium mEq/L 1.3 10.7 4.2 0.80 Blood 79857
-RDW % 10.6 33.8 13.8 2.19 Blood 82266
-Red Blood Cells m/uL 0.65 8.06 4.25 0.96 Blood 82266
-Sodium mEq/L 88.0 179.0 139.0 5.0 Blood 79865
-Troponin T ng/mL 0.0 17.36 0.06 0.13 Blood 41806
-Urea Nitrogen mg/dL 1.0 263.0 18.0 13.0 Blood 82522
-White Blood Cells K/uL 0.0 632.1 8.1 4.60 Blood 82367
-pCO2 mm Hg 8.0 189.0 44.0 17.0 Blood 8008
-pH units 5.0 9.0 6.0 1.0 Urine 20968
-Table 8: Summary of laboratory values across samples considered in this study, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records for each parameter.
-Table 8 provides details on the laboratory values considered in this work, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records.
-During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence.
-The specific thresholds for outliers were as follows: absolute basophil count above 20, absolute eosinophil count above 20, absolute lymphocyte count above 100, alanine aminotransferase above 2000, alkaline phosphatase above 2000, aspartate aminotransferase above 2000, creatine kinase above 2000, glucose above 2000, lactate above 2000, and platelet count above 2000.
-Laboratory values were extracted from the laboratory events table from MIMIC-IV, where for each record we matched all the patient records of the desired tests available between patient record in-time and the 1.5 hours window.
-Appendix B Models
-Hyperparameter Value
-Block of layers 4
-S4 model copies 512
-S4 state size 8
-Optimizer Adam
-Learning rate 0.001
-Weight decay 0.001
-learning rate schedule constant
-Batch size 64
-Epochs 20
-Table 9: S4 model hyperparameters
-Table 9 outlines the hyperparameters employed in the S4 model.
-The architecture consists of four blocks of layers, with each block containing 512 copies of the S4 model.
-The state size within the S4 model is set to 8.
-For optimization, the AdamW optimizer is utilized with a learning rate and weight decay both set to 0.001.
-The learning rate schedule is maintained constant throughout training.
-A batch size of 64 samples is used for each training iteration, spanning a total of 20 epochs.
-The training objective is to minimize the binary cross-entropy loss.
-During training, we apply a model selection strategy on the best performance (AUROC) on the validation set which usually converges earlier than the final number of epochs.
-As previously described in Appendix A, with the goal of not discarding samples whose any of the deterioration targets happens before the 1.5-hour window or the target event cannot be clearly defined, we introduce a special token in the label space (-999) for which we account during and training and evaluation to be discarded.
-As of XGboost models, we fit these models with the default hyperparameter values in a binary setting for the later AUROC aggregation across settings, models which in principle do not intrinsically allow the inclusion of samples during training with the special token to be discarded.

### v4

@@
-a Benchmark Dataset for Diagnoses and Deterioration Prediction in Emergency Medicine
+Enhancing clinical decision support with physiological waveforms – a multimodal benchmark in emergency care
 Abstract
-Background: A clinically meaningful comparative assessment of medical decision support in emergency care is challenging due to a lack of appropriate datasets with multimodal input modalities and comprehensive prediction task.
-This hampers measurable progress in the field.
-Results: We introduce a dataset based on MIMIC-IV, a benchmarking protocol, and initial results for evaluating multimodal decision support in the emergency department (ED).
-We use diverse data modalities from the first 1.5 hours after patient arrival, including demographics, biometrics, vital signs, lab values, and electrocardiogram waveforms.
-We analyze 1443 clinical labels across two contexts: predicting diagnoses and patient deterioration.
-Our diagnostic model achieves an AUROC score over 0.8 in a statistically significant manner for 609 out of 1428 conditions, including cardiac conditions like myocardial infarction and non-cardiac conditions such as renal disease and diabetes.
-The deterioration model scores above 0.8 in a statistically significant manner for 14 out of 15 targets, including critical events like cardiac arrest, mechanical ventilation, intensive care unit admission, as well as short- and long-term mortality.
-Furthermore, we provide one of the first robust demonstrations of the significant impact of raw waveform input data on model performance.
-Conclusions: This study highlights the proposed dataset as a unique resource to foster progress towards measurable progress in the domain of algorithmic decision support in emergency care.
-The presented multimodal baseline models showcase the potential of diagnostic decision support in the field and provide strong incentives for including raw waveform data.
+Background: AI-driven prediction algorithms have the potential to enhance emergency medicine by enabling rapid and accurate decision-making regarding patient status and potential deterioration.
+However, the integration of multimodal data, including raw waveform signals, remains underexplored in clinical decision support.
+Methods: We present a dataset and benchmarking protocol designed to advance multimodal decision support in emergency care.
+Our models utilize demographics, biometrics, vital signs, laboratory values, and electrocardiogram (ECG) waveforms as inputs to predict both discharge diagnoses and patient deterioration.
+Results: The diagnostic model achieves area under the receiver operating curve (AUROC) scores above 0.8 for 609 out of 1,428 conditions, covering both cardiac (e.g., myocardial infarction) and non-cardiac (e.g., renal disease, diabetes) diagnoses.
+The deterioration model attains AUROC scores above 0.8 for 14 out of 15 targets, accurately predicting critical events such as cardiac arrest, mechanical ventilation, ICU admission, and mortality.
+Conclusions: Our study highlights the positive impact of incorporating raw waveform data into decision support models, improving predictive performance.
+By introducing a unique, publicly available dataset and baseline models, we provide a foundation for measurable progress in AI-driven decision support for emergency care.
 keywords:
-Medical Decision Support; Multimodal Data; Emergency Department; Diagnostic Prediction; Patient Deterioration.
-journal: general
-{keypoints*}
-  • Introduction of a novel dataset based on MIMIC-IV for benchmarking multimodal decision support in the emergency department.
-  • Evaluation of 1443 clinical labels for diagnostic and patient deterioration predictions, achieving high AUROC scores.
-  • Demonstration of significant performance improvements by incorporating raw ECG waveform data.
-  • Highlighting the dataset’s comprehensive feature set collected within the first 1.5 hours after patient arrival.
-  • Potential impact on improving decision-making processes in acute and emergency medicine.
+Deep-learning , Emergency Department , Machine learning , Medical Decision Support , Multimodal Data , Patient Diagnostics , Patient Deterioration.
+journal: Computers in Biology and Medicine
+url]https://uol.de/en/ai4health
+[…]
+[label1]organization=AI4Health Division, Carl von Ossietzky Universität Oldenburg, addressline=Ammerländer Heerstraße 114-118, city=Oldenburg, postcode=26129, state=Lower Saxony, country=Germany
+[…]
+[label2]organization=Department of Internal Medicine, Department of Acute Care, and Department of Clinical Pharmacy & Pharmacology, University Medical Center Groningen, addressline=Hanzeplein 1, city=Groningen, postcode=9713, state=Groningen, country=Netherlands
 1 Introduction
 Artificial intelligence (AI) is an emerging field that has significantly enhanced healthcare and medicine [1] in areas such as precision medicine [2] and drug discovery [3].
-The implementation of AI models is particularly relevant in acute and emergency medicine, where clinicians must address critical conditions within short time frames to make optimal clinical decisions.
+The implementation of AI models is particularly relevant in acute and emergency care, where clinicians must address critical conditions within short time frames to make optimal clinical decisions.
 AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8, 9].
@@
 However, many of these studies have limitations, such as predicting only a limited set of specific conditions [10, 4], offering short time horizons for survival predictions, and requiring costly diagnostic tests [11].
+Despite their potential, current state-of-the-art AI models for forecasting specific medical conditions rely primarily on standard clinical features (e.g., demographics, laboratory results) and neglect raw physiological signals such as waveforms.
+These signals are known to be highly informative for a diverse set of diseases [12] and represent one example of raw sensor data.
+They hold the promise of providing complementary information to standard clinical features to enhance the accuracy and robustness of prediction models.
+The prediction models used vary widely, as noted in a recent meta-review [13] in the context of sepsis prediction, but typically cannot be compared directly due to differences in experimental setup.
+This highlights the need for structured benchmarking protocols.
+Unlike mere evaluation, which often focuses on fixed metrics for a single study, benchmarking establishes common datasets, tasks, and performance baselines, ensuring that advancements in AI-driven clinical decision support are measurable, reproducible, and comparable across studies.
+Benchmarking, as one component of standardized performance evaluation, also plays a crucial role in validation and certification processes.
+These processes represent a significant hurdle that decision support systems must overcome to reach clinical practice.
 Healthcare datasets are essential for advancing medical research and innovation.
@@
 Nonetheless, these datasets still have limitations, including size constraints [14], lack of open-source availability [18, 19], and narrow scopes in terms of prediction tasks [14, 15, 19, 16, 18, 17].
-Benchmarking is crucial for evaluating the performance and reliability of novel AI-models, ensuring their applicability and robustness in real-world clinical settings.
-In this study, we investigate multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities collected within the first 1.5 hours of patient arrival at the ED department such as demographics, biometrics, vital parameters trends, laboratory trends, and electrocardiogram (ECG) waveforms, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
-2 Analyses
+In this study, we investigate a multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities collected within the first 1.5 hours of patient arrival at the ED department such as demographics, biometrics, vital parameters trends, laboratory trends, and electrocardiogram (ECG) waveforms, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
+Compared to approaches using a single data modality, which may offer a limited view of the patient’s condition, we demonstrate that fusing multiple data sources is critical for improving decision-making accuracy in the complex environment of emergency departments.
+This multimodal approach is essential for predicting immediate as well as long-term outcomes, enabling more comprehensive and accurate assessments of patient status.
+To summarize, we put forward the following technical contributions:
+  1. 1.
+    We propose the MDS-ED dataset along with a corresponding benchmarking protocol which stands out from literature approaches as detailed in Section 2.2.
+  2. 2.
+    We put forward strong unimodular baseline models based on best-practices from the literature, see Section 2.3.
+  3. 3.
+    We propose powerful multimodal prediction models that outperform these strong unimodal baselines in all the scenarios considered, as demonstrated in Section 3.
+To our knowledge, this is the first robust demonstration of the added value of incorporating raw waveforms into clinical prediction models in an acute care setting.
+2 Methods
+2.1 Clinical workflow and dataset creation
+Figure 1: Pipeline outlines the MDS-ED clinical workflow, which involves feature collection encompassing patient demographics, biometrics (such as height, weight, and BMI), vital parameters and trends, laboratory values and trends, and ECG waveform data to address two clinically relevant prediction scenarios: predicting patient discharge diagnoses out of 1428 cardiac and non-cardiac ICD10-CM codes and predicting patient deterioration according to 15 clinical deterioration measures.
+Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay.
+MDS-ED was created by linking ECG waveforms MIMIC database, specifically, MIMIC-IV-ECG [20] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [21].
+The MIMIC dataset was chosen for its extensive collection of patient records across diverse demographic groups, encompassing thousands of features from clinical notes, lab results, vital signs, and full waveform data.
+This multimodal breadth enhances AI models predictive performance and clinical decision support.
+Figure 2: Schematic representation summarizing the creation process of the MDS-ED dataset [22] underlying this work.
+The process starts from four different source datasets (MIMIC-IV-ECG, MIMIC-IV, MIMIC-IV-ED, and MIMIC-IV-ECG-ICD) from which we select patients aged 18 years or older where an ECG was collected within the first 90 minutes of arrival at the ED department.
+Different target values collected at different intervals.
+On the resulting samples, we procedded to apply a features outlier removal which is primarily error-based, excluding unrealistic values, never-registered extremes, or negative values when the minimum is zero, see Appendix C and [22] for details.
+Similarly, an imputation mask creation in which we apply median imputation from the train set to validation and test, adding binary masks to indicate imputed values, helping the model learn missingness patterns.
+Finally, we capture trends through engineered features such as summary statistics (mean, median, min, max, standard deviation), first and last values, rate of change, and slope of fitted linear model based on values within the first 90 minutes after arrival.
+The resulting MDS-ED dataset comprises 1428 diagnostic labels and 15 deteriorations labels across 129,057 samples from 71,098 patients collected from 121,195 unique visits.
+The input features cover a single 10s, 12-lead ECG in addition to 470 tabular features (excluding binary masking columns).
+The MDS-ED dataset was constructed by integrating multiple sources to create a comprehensive resource for emergency medicine research.
+This process involved merging ECG recordings with diverse patient characteristics and clinical outcomes.
+Specifically, diagnostic ICD-10 codes and stratified splits were obtained from MIMIC-IV-ECG-ICD [12], while ECG recordings were sourced from MIMIC-IV-ECG [20].
+Clinical data were extracted from MIMIC-IV [21], utilizing tables related to patient admissions, diagnoses, laboratory tests, ICU stays, and procedures.
+Additionally, emergency department-specific information, including stays, diagnoses, medication administration, vital signs, and triage details, was gathered from MIMIC-IV-ED [21].
+The dataset creation process in visualized schematically in Figure 2.
+A complete description of the MDS-ED dataset along with the full preprocessing code for its creation is available on PhysioNet[22].
+Prediction tasks and targets
+In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration.
+For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [12], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG).
+To this end, we frame the task as a multilabel classification task in terms of international classification of diseases clinical modification 10 (ICD-10 CM) codes.
+As in [12], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits, and propagate codes to parent codes up to the third digit.
+In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions.
+For patient deterioration, we build on a consensus definition [23] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA).
+Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days.
+While ordinary predictive models often focus on immediate deterioration, our approach extends beyond standard scenarios by considering a broader temporal scope and a more comprehensive set of clinical events.
+Long-term prediction tasks, such as mortality at extended time horizons, pose unique challenges due to the evolving nature of patient conditions, while discharge diagnoses rely on accumulated clinical knowledge throughout the hospital stay and could be highly influential if identified early.
+By addressing both immediate and long-term risks, our approach ensures a more holistic and clinically relevant decision support framework.
+With the two main groups of diagnoses and deterioration, we aim to cover a wide spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels.
+See Table 7 in the supplementary material for a detailed discussion of the prediction targets.
+Features
+While aiming to include mainly data at triage, we also include irregularly sampled vital parameters and lab values captured during the 1.5h data collection window.
+To capture trends, we compute simple sample-wise statistical aggregation functions per vital parameter and laboratory values, namely mean, median, minimum, maximum, standard deviation, first and last values, rate of change of them between first and last value, as well as the slope of a linear model fitted on the minutes difference between value collection and arrival as independent variable and the actual values as dependent variables.
+We follow [24] to select an appropriate set of laboratory values that avoids a bias towards particular prediction targets.
+We obtain a total of 470 features of clinical routine data across all feature modalities in a tabular format plus a 12-lead 10-second ECG waveform per sample.
+Similarly, the ECG feature set is composed of RR interval, P-onset, QRS-onset, QRS-end, and T-end in milliseconds (ms), whereas P-axis, QRS-axis, and T-axis in degrees.
+Train-test splits
+MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,057 samples.
+We leverage the stratified splits provided along with the dataset accompanying [12], which includes stratification based on gender, age bins, and discharge diagnoses.
+We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1.
+2.2 Related work
+Table 1: Direct comparison with related works in terms of dataset size, features, availability, and number of target labels.
+Group Detail MIMIC-ED-Assist VitalML MEME EHRSHOT MC-BEC MDS-ED
+Source [14] [15] [19] [16] [18] This work
+Population ED ED Longitudinal Longitudinal ED ED
+Size Patients 25714 N/A 947028 6739 63389 71098
+Visits 32356 19847 N/A 921499 102731 121195
+Features Demographics
+Biometrics
+Vital parameters ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$)
+Lab. values ✓($\mathbb{T}$) ✓($\mathbb{T}$)
+Waveforms ✓($\mathbb{E}$)
+Chief complaint
+Medications
+Tasks Labels 3 3 4 15 7 1443
+Availability Open source
+  • 1.
+    We use diverse symbology to express the contribution where ✓ = available, ✗ = unavailable, $\mathbb{E}$ = available in the form of embeddings, and $\mathbb{T}$ = available in the form of trends or at least two sampled values.
+Table 1 contains a direct comparison with related works in terms of population, dataset size, features, availability, and number of target labels.
+[14] proposed the MIMIC-ED-Assist benchmark to reduce ED length of stay by flagging high-risk patients using triage features.
+[15] proposed VitalML to predict deterioration in the next 90 minutes from the first 15 minutes of monitoring.
+[19] proposed a multiple embedding model for EHR data (MEME) to generate pseudo-notes from tabular data and predict ED disposition, discharge location, ICU admission, and inpatient mortality utilizing a natural language processing approach.
+[16] proposed a few shot foundational model EHRSHOT approach to predict operational outcomes, developed diagnoses, as well as anticipate lab test results and image findings.
+[18] proposed MC-BEC to predict patient deterioration, disposition, and revisit from features collected including diagnoses from the first 15 minutes after the patient is assigned to a room.
+Overall, our proposed MDS-ED main contributions in terms of the dataset are: 1) Comprehensive size: MDS-ED is situated in the first position regarding a large number of patients and second in the number of visits in the open-source domain, despite restricting towards the investigated setting of the first 1.5 hours of ED arrival.
+2) Diverse input features: MDS-ED is situated in the first position in terms of feature modalities in the open-source domain as it incorporates a comprehensive set of features, including demographics, biometrics, vital parameter trends, laboratory values trends, and ECG waveforms.
+This is more extensive than most compared datasets.
+Intentionally, we decided to exclude chief complaints as their unstructured nature presents a more challenging step towards external validation with the addition of more variance, see [19].
+Similarly, we exclude previous patient medications as they involve irregularities in administration times, and dosages, as well as potential bias towards task prediction, e.g., certain medications are given for specific gravity and stages of diagnoses.
+3) Comprehensive prediction targets: MDS-ED proposes 1,443 target labels, significantly more than other datasets, which typically cover fewer and too narrow in scope tasks.
+4) Accessibility: Similar to [14, 15, 16], MDS-ED is open source, encouraging further research and collaboration.
+2.3 Model architectures
+To propose an initial benchmark, we experiment with diverse settings based on data modalities input such as (1) clinical routine data, (2) ECG features only, (3) ECG waveforms only, (4) ECG features + clinical routine data, and (4) ECG waveforms + clinical routine data.
+For waveforms (3) and (5), inspired by recent successful applications in the field of physiological time series [12, 25, 26] we employed an S4 classifier based on structured state space models with four layers [27] (relying on hyperparameter choices from prior work[26], i.e, using state dimension 8 and model dimension 512), which has been shown to outperform more conventional deep learning architectures, such as modern convolutional neural networks, recurrent neural networks or transformer models, as signal encoder [12, 25, 26].
+For the multimodal models in (5), we use a 3-layer MLP as tabular encoder, where the three categorical features (gender, race, acuity score) are first processed through appropriate embedding layers.
+The outputs of the signal encoder and the tabular encoder are then fused through concatenation.
+We apply median imputation using training set statistics.
+In addition, we introduce a binary column per input variable to indicate imputed values.
+These binary columns are provided as additional inputs next to the original tabular features.
+In A in the supplementary material, we present an ablation study comparing different imputation strategies.
+A systematic exploration of other tabular encoder models or more complex multimodal fusion schemes beyond the basic choices from above is deferred to future work.
+For (1), (2) and (4) we fit extreme gradient boosting (XGBoost) decision trees, which widely accepted state-of-the-art for prediction models on tabular data.
+We used standard hyperparameter defaults commonly applied in gradient-boosting methods.
+Additional details on model architectures and model hyperparameters can be found in D Table 11.
+2.4 Training and evaluation
+For the S4 models, we used AdamW [28] as the optimizer, with both learning rate and weight decay set to 0.001, and maintained a constant learning rate schedule.
+The training was conducted with a batch size of 64 samples over 20 epochs The hyperparameter settings coincide with those from an earlier unimodal investigation [12], which underscores the robustness of our findings without excessive hyperparameter tuning.
+Overfitting was mitigated through model selection based on validation performance, making the total number of epochs irrelevant as long as it was sufficiently large for convergence.
+The training objective was to minimize binary cross-entropy loss.
+For all the models, we evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC).
+To assess statistical uncertainty resulting from the test set’s finite size and specific composition, we employ empirical bootstrap on the test set with $n=1000$ iterations.
+We report 95% confidence intervals for both macro AUROC and individual label AUROCs.
+Given [12] evaluation protocol, we use only the first record per visit per patient to avoid bias.
+We use macro AUROC as the primary metric due to its status as most widely used ranking-based metric.
+It characterize a model’s overall discriminative power without the necessity to fix custom decision thresholds.
+The suitability of AUROC compared to the other commonly used metric in the presence of label imbalance, the area under the precision recall curve (AUPRC), was confirmed in a recent theoretical and empirical work [29].
+Additional details of training and evaluation procedures can be found in D.
+3 Results
 3.1 Benchmarking predictive performance
@@
 In the following discussion, we restrict ourselves to the discussion of the most comprehensive multimodal model based on ECG waveforms and clinical routine data.
+In B, we present a more finegrained comparison across different ICD10 chapters and specific deterioration targets, where the multimodal deep learning consistently outperforms unimodal baseline models.
 3.2 Task-dependent predictive performance
@@
 In total, the model predicted 609 out of 1,428 individual ICD diagnoses with high accuracy, which we define in this work as conditions where the lower bound of the 95% bootstrap confidence interval for AUROC exceeds 0.80.
+A more fine-grained comparison of model performance across different input settings in the deterioration prediction task reveals notable trends in Table 5, where the inclusion of ECG waveforms alongside clinical routine data consistently improves performance over ECG features and clinical routine data, with relative improvements ranging from 2.36% to 13.06%.
+The highest gains are observed for XII (Skin, 13.06%), VI (Nervous, 10.75%), and XIX (Injury, Poisoning, 9.95%), suggesting that continuous waveform signals provide critical insights over features into these non-cardiac groups.
+In contrast, smaller improvements are seen for III (Blood, 2.36%) and IV (Endocrine, 5.25%), indicating that static ECG features might already capture most of the predictive signal in these cases.
 Table 4: Detailed breakdown of model performance in the deterioration category.
@@
 Overall, for clinical deterioration, the model achieves an AUROC of 0.9070, whereas, for ICU admissions, the model reports an overall AUROC of 0.9063, and for mortality predictions also exhibits high performance, with an overall AUROC of 0.9168.
+In Tables 5 and 6 in the supplementary material, we present a more fine-grained comparison of model performance across different input settings for the deterioration prediction task.
+As in the diagnostic case, the multimodal model leveraging waveforms outperforms the corresponding model leveraging only ECG features in addition to clinical routine data in all cases, most notably in the clinical deterioration category with a relative improvement of more than 9%.
+On the most finegrained level in Table 6, one observes relative improvements ranging from 1.41% (365-day mortality) to 38.08% (severe hypoxemia).
+The most substantial gains occur for severe hypoxemia (38.08%) and inotrope administration (16.57%), suggesting that waveform-based inputs contribute significantly to an early identification of respiratory failure and cardiac dysfunction.
+Conversely, for longer-term mortality outcomes (e.g., 90-day to 365-day), the added benefit of waveform data diminishes, as patient trajectories become increasingly influenced by non-acute factors that are already reliably captured by clinical routine features and ECG features.
 4 Discussion
@@
 To the best of our knowledge, this is the first statistically robust demonstration of the added value of raw ECG waveform input against ECG features for clinically relevant prediction tasks such as diagnoses and deterioration prediction.
-Thirdly, for the unimodal models, in the deterioration task, the clinical routine data model outperforms ECG features only and ECG waveforms only, however, in the diagnoses task, the ECG waveforms only outperforms the other 2 settings, we hypothesize that for the deterioration task, the clinical routine data apart of including a rich set of clinical features (demographics, biometrics, vital parameters trends, and laboratory values trends) against only an single ECG either in features or waveform, it also includes trends over time which aligns with the task definition of deterioration.
+Thirdly, for the unimodal models, in the deterioration task, the clinical routine data model outperforms ECG features only and ECG waveforms only, however, in the diagnoses task, the ECG waveforms only outperforms the other two settings, we hypothesize that for the deterioration task, the clinical routine data apart of including a rich set of clinical features (demographics, biometrics, vital parameters trends, and laboratory values trends) against only an single ECG either in features or waveform, it also includes trends over time which aligns with the task definition of deterioration.
 Despite this, we believe that a single ECG snapshot can achieve high performances for both tasks, but also we believe that the inclusion of multiple ECGs over time instead of just a single snapshot would allow us to capture more meaningful deterioration and potentially diagnoses trends.
+Fourthly, direct comparison to literature results is challenging due to the lack of publicly available datasets and the corresponding scarcity of models that integrate both clinical routine data and waveform data.
+Among existing studies, [18] is one of the few incorporating waveform-based models, but differences in dataset construction and task formulation limit direct comparisons.
+Consequently, from a methodological perspective, we evaluate our approach using state-of-the-art models, including an ensemble tree-based model for tabular data and a signal encoder with proven efficacy, as previously presented in [12].
 4.2 Clinical significance
@@
 Despite the promising results, there are several limitations to consider.
-Firstly, the presented models rely on the specifics of Beth Israel Deaconess Medical Center in terms of patient demographics, available resources, process characteristics, and data collection practices.
-We carefully quantify the statistical uncertainties of our results and are convinced of the robustness of our results for this particular scenario within the limits of a retrospective evaluation.
-However, we cannot make any statements about the transferability of our findings to other hospitals or deployment scenarios as this would require an external validation dataset, see [19] for a recent perspective on the role of external validation for clinical decision support.
-Secondly, the reliance on data collected within the first 1.5 hours of the ED visits may limit the model’s applicability.
+Firstly, diagnostic labels were inferred from ICD10 codes, which are subject to biases and incompleteness as they are primarily used for billing purposes and are not assigned purely based on clinical considerations.
+Clinical deterioration labels are comparably reliable as they are sourced from the same hospital system.
+Post-discharge mortality might again be subject to inaccuracies as it originates from a different source system.
+Secondly, the presented models rely on the specifics of Beth Israel Deaconess Medical Center in terms of patient demographics, available resources, process characteristics, and data collection practices.
+The presented results allow statements about the generalizability of the proposed models to unseen patients with well-defined statistical uncertainty measures, as far as this is possible.
+However, it does not allow any statements about the transferability of these findings to other hospitals.
+It is important to note that generalization to other sites is challenging not only due to different patient demographics but also due to available resources, process characteristics, and data collection practices.
+This extends as far as certain labels, such as ICU admission, that are directly affected by process characteristics.
+The challenges of external validation in the context of complex clinical decision support systems are nicely summarized in a recent article [30].
+Thirdly, all considered prediction scenarios rely on data collected within the first 1.5 hours of the ED visits, which may limit the model’s applicability.
 We deliberately decided to focus on this setting to keep the dataset homogeneous.
-However, there is also a clear clinical case for a prediction model only incorporating information available during triage.
-A similar dataset to the one proposed could be created, but this task is left for future research.
-Thirdly, while we focus on the construction of a benchmark dataset and provide the first results for it, we do not provide any insights into the trained models beyond a discussion of the model performance for specific conditions.
-This could be achieved for example through attribution methods proposed in the context of explainable AI, however, mostly in uni-modular settings [20], to verify that the models do not exploit spurious correlations to reach their strong performance.
-Again, a detailed investigation also of causal vs.
-associational attributions along the lines of [21] is beyond the scope of this submission.
-Finally, although the open-source dataset fosters innovation, the implementation of a large-scale single end-to-end model with a wide range of target variables such as the diagnoses case might result in challenges during deployment in resource-constrained hardware, see 9 in the supplementary material for a more extensive discussion.
+There are two directions that are worth exploring in future work.
+On the one hand, there is a clear clinical case for a prediction model only incorporating information available during triage.
+On the other hand, it might be instructive to extend the time horizon beyond 1.5 hours in particular to improve the performance for deterioration prediction tasks.
+In both cases, datasets similar to the one proposed could be created, but this task is left for future research.
+Finally, structured benchmarks through open-source datasets is only a first step towards measurable progress in the domain.
+Translating these research findings into clinical applications poses significant additional challenges.
+For example, deploying large-scale models will require substantial investments in hardware, in particular for real-time applications.
+Clinical decision support systems have to be carefully validated on diverse patient population and have to undergo certification, have to be properly integrated into clinical workflows and have to be closely monitored after deployment to detect early signs of performance deterioration due to time-dependent dataset shifts.
+To ensure clinical utility, future work should prioritize validation across diverse patient populations and healthcare settings.
+While many AI algorithms are developed and published, only a small fraction undergo prospective validation or certification for real-world deployment.
+Regulatory approval and integration into existing clinical workflows present significant challenges, which may impact clinician trust and adoption.
+A structured benchmarking process could help to streamline this by ensuring a more standardized performance evaluation.
 4.4 Future research directions
-The proposed dataset and benchmark protocol build on input parameters that are typically available in clinical workflows and address a set of comprehensive, clinically meaningful prediction tasks.
+The proposed dataset and benchmark protocol are based on input parameters that are typically available in clinical workflows and address a set of comprehensive and clinically meaningful prediction tasks.
 We therefore envision it to represent a meaningful baseline to benchmark prediction algorithms in the field.
-An obvious extension would be to include free-text data along the lines of [16].
+In addition to addressing the limitations mentioned in the previous sections, we consider pursuing the following technical research directions as particularly promising:
+Future research directions could focus on expanding the scope of input variables beyond the current feature set, incorporating additional clinical and physiological signals to enhance predictive performance.
+An obvious extension along these lines would be to include free-text data as in the recent of Hager et al [17].
+This could enhance the predictive performance by capturing richer clinical context, uncovering implicit relationships, and incorporating nuanced information that structured data alone might miss.
 Going beyond MIMIC-IV, models developed on the proposed dataset could foreshadow the development of more comprehensive models on more specialized datasets such as [31], for example including molecular data as input and/or prediction targets, to address more fine-grained diagnostic or deterioration prediction targets.
+Another promising avenue involves using multiple waveform snapshots over time for deterioration prediction, rather than relying solely on a single observation, thereby capturing dynamic changes in patient conditions more effectively.
+Finally, our work naturally focuses on quantitative accuracy.
+However, there are many quality dimensions beyond quantitative accuracy that are important for clinical applications, such as explainability, different aspects of robustness, uncertainty quantification or privacy.
+As a specific example, it would be interesting to investigate whether specific ECG patterns contribute to the performance gains observed when using full ECG waveforms instead of extracted features.
+Related methods for explaining ECG-based predictions have been explored in [32], and further work could extend these approaches to assess causal versus associational attributions along the lines of [33].
+As a second specific example, even though we work with de-identified structured data in this work, it is important to develop models in line with privacy requirements.
+Here, emerging privacy-preserving techniques, as detailed for example in [34], will become essential to ensure compliance in clinical decision support systems, especially with the potential future inclusion of free-text data.
 5 Potential implications
@@
 However, questions of algorithmic fairness remain to be investigated in detail in future work.
-5 Methods
-5.1 Clinical workflow and dataset creation
-Figure 1: Pipeline outlines the MDS-ED workflow, which involves feature collection encompassing patient demographics, biometrics (such as height, weight, and BMI), vital parameters and trends, laboratory values and trends, and ECG waveform data to address two clinically relevant prediction scenarios: predicting patient discharge diagnoses out of 1428 cardiac and non-cardiac ICD10-CM codes and predicting patient deterioration according to 15 clinical deterioration measures.
-Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay.
-MDS-ED was created by linking ECG waveforms from the medical information mart for intensive care (MIMIC) database, specifically, MIMIC-IV-ECG [23] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [24].
-Prediction tasks and targets
-In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration.
-For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [12], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG).
-To this end, we frame the task as a multilabel classification task in terms of international classification of diseases clinical modification 10 (ICD-10 CM) codes.
-As in [12], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits, and propagate codes to parent codes up to the third digit.
-In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions.
-For patient deterioration, we build on a consensus definition [25] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA).
-Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days.
-With these two main groups, we aim to cover a wide range spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels, see Table 5 in the supplementary material for a detailed target definition discussion.
-Features
-While aiming to include mainly data at triage, we also include irregularly sampled vital parameters and lab values captured during the 1.5h data collection window.
-To capture trends, we compute simple sample-wise statistical aggregation functions per vital parameter and laboratory values, namely mean, median, minimum, maximum, standard deviation, first and last values, rate of change of them between first and last value, as well as the slope of a linear model fitted on the minutes difference between value collection and arrival as independent variable and the actual values as dependent variables.
-We follow [26] to select an appropriate set of laboratory values that avoids a bias towards particular prediction targets.
-We obtain a total of 470 features of clinical routine data across all feature modalities in a tabular format plus a 12-lead 10-second ECG waveform per sample.
-Similarly, the ECG feature set is composed of RR interval, P-onset, QRS-onset, QRS-end, and T-end in milliseconds (ms), whereas P-axis, QRS-axis, and T-axis in degrees.
-MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,057 samples.
-We leverage the stratified splits provided along with the dataset accompanying [12], which includes stratification based on gender, age bins, and discharge diagnoses.
-We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1.
-For deep learning models with clinical routine data, we apply a median imputation learned from train set to validation and test for each column and we apply a masking strategy in which we add an additional binary column to indicate which values were imputed to the model.
-See the supplementary material for further preprocessing and definitions details as well as descriptive statistics on demographics, biometrics, vital parameters and laboratory values in Tables 6-8.
-5.2 Related work
-Table 4: Direct comparison with related works in terms of dataset size, features, availability, and number of target labels.
-Group Detail MIMIC-ED-Assist VitalML MEME EHRSHOT MC-BEC MDS-ED
-Source [13] [14] [18] [15] [17] This work
-Population ED ED Longitudinal Longitudinal ED ED
-Size Patients 25714 N/A 947028 6739 63389 71098
-Visits 32356 19847 N/A 921499 102731 121195
-Features Demographics
-Biometrics
-Vital parameters ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$)
-Lab. values ✓($\mathbb{T}$) ✓($\mathbb{T}$)
-Waveforms ✓($\mathbb{E}$)
-Chief complaint
-Medications
-Tasks Labels 3 3 4 15 7 1443
-Availability Open source
-{tablenotes}
-We use diverse symbology to express the contribution where ✓ = available, ✗ = unavailable, $\mathbb{E}$ = available in the form of embeddings, and $\mathbb{T}$ = available in the form of trends or at least two sampled values.
-Table 4 contains a direct comparison with related works in terms of population, dataset size, features, availability, and number of target labels.
-[13] proposed the MIMIC-ED-Assist benchmark to reduce ED length of stay by flagging high-risk patients using triage features.
-[14] proposed VitalML to predict deterioration in the next 90 minutes from the first 15 minutes of monitoring.
-[18] proposed a multiple embedding model for electronic health records (EHR) (MEME) to generate pseudo-notes from tabular data and predict ED disposition, discharge location, ICU admission, and inpatient mortality utilizing a natural language processing approach.
-[15] proposed a few shot foundational model EHRSHOT approach to predict operational outcomes, developed diagnoses, as well as anticipate lab test results and image findings.
-[17] proposed MC-BEC to predict patient deterioration, disposition, and revisit from features collected including diagnoses from the first 15 minutes after the patient is assigned to a room.
-Overall, our proposed MDS-ED main contributions in terms of the dataset are: 1) Comprehensive size: MDS-ED is situated in the first position regarding a large number of patients and second in the number of visits in the open-source domain, despite restricting towards the investigated setting of the first 1.5 hours of ED arrival.
-2) Diverse input features: MDS-ED is situated in the first position in terms of feature modalities in the open-source domain as it incorporates a comprehensive set of features, including demographics, biometrics, vital parameter trends, laboratory values trends, and ECG waveforms.
-This is more extensive than most compared datasets.
-Intentionally, we decided to exclude chief complaints as their unstructured nature presents a more challenging step towards external validation with the addition of more variance, see [18].
-Similarly, we exclude previous patient medications as they involve irregularities in administration times, and dosages, as well as potential bias towards task prediction, e.g., certain medications are given for specific gravity and stages of diagnoses.
-3) Comprehensive prediction targets: MDS-ED proposes 1,443 target labels, significantly more than other datasets, which typically cover fewer and too narrow in scope tasks.
-4) Accessibility: Similar to [13, 14, 15], MDS-ED is open source, encouraging further research and collaboration.
-5.3 Model architectures
-To propose an initial benchmark, we experiment with diverse settings based on data modalities input such as (1) clinical routine data, (2) ECG features only, (3) ECG waveforms only, (4) ECG features + clinical routine data, and (4) ECG waveforms + clinical routine data.
-For waveforms (3) and (5), inspired by recent successful applications in the field of physiological time series [12, 27, 28] we employed an S4 classifier based on structured state space models with four layers [29] (with state dimension 8 and model dimension 512 as in [28]).
-For the multimodal models in (5), we use a 3-layer MLP as tabular encoder, where the three categorical features (gender, race, acuity score) are first processed through appropriate embedding layers.
-The outputs of the signal encoder and the tabular encoder are then fused through concatenation.
-For (1), (2) and (4) we fit extreme gradient boosting (XGBoost) decision trees.
-Additional details on model architectures and model hyperparameters can be found in 9 Table 9.
-5.4 Training and evaluation
-For the S4 models, we used Adam Weight Decay Optimization (AdamW) as the optimizer, with both learning rate and weight decay set to 0.001, and maintained a constant learning rate schedule.
-The training was conducted with a batch size of 64 samples over 20 epochs which usually converged earlier and model selection on the validation set.
-The training objective was to minimize binary cross-entropy loss.
-For all the models, we evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC).
-To assess statistical uncertainty resulting from the test set’s finite size and specific composition, we employ empirical bootstrap on the test set with $n=1000$ iterations.
-We report 95% confidence intervals for both macro AUROC and individual label AUROCs.
-Given [12] evaluation protocol, we use only the first record per visit per patient to avoid bias.
-Additional details of training and evaluation procedures can be found in 9.
-6 Availability of source code and requirements
+5.1 Data and code availability
+  • 1.
     Project name: MDS-ED
+  • 2.
     Project home page: [22].
-  • Project code repository: [31].
+  • 3.
+    Project code repository: https://github.com/AI4HealthUOL/MDS-ED
+  • 4.
     Operating system(s): Platform independent
+  • 5.
     Programming language: Python
+  • 6.
     Other requirements: Python 3.10.8, PyTorch 1.13.0, and PyTorch Lightening 1.8.0.
-  • License: MIT
-7 Declarations
-7.1 List of abbreviations
-AdamW: Adam Weight Decay Optimization, AI: Artificial Intelligence, AUROC: Area Under the Receiver Operating Characteristic Curve, ECG: Electrocardiogram, ECMO: Extracorporeal Membrane Oxygenation, ED: Emergency Department, EHR: Electronic Health Records, ICD CM: International Classification of Diseases Clinical Modification, ICU: Intensive Care Unit, IHCA: In-Hospital Cardiac Arrest, MDS-ED: Multimodal Decision Support for Emergency Department, MIMIC: Medical Information Mart for Intensive Care, MLP: multi-layer perceptron, XGBoost: Extreme Gradient Boosting.
-7.2 Consent for publication
-Not applicable
+5.2 Author contributions
+JMLA and NS were responsible for the conceptualization of the project.
+Data curation was performed by JMLA and NS.
+JMLA and NS implemented the prediction models, with JMLA conducting the formal analysis under NS’s supervision.
+HB provided critical revisions for clinical content.
+JMLA produced the original draft, and NS and HB contributed to the review and editing.
+All authors participated in the investigation, methodology, and validation, and they approved the final version for publication.
+Computing resources were provided through Carl von Ossietzky University of Oldenburg.
+5.3 Generative AI statement
+Generative AI was used for grammar correction but not for text generation in this manuscript.
 5.4 Competing interests
 The authors have declared no competing interests.
-7.4 Author’s Contributions
-JMLA and NS conceived and designed the project.
-NS and JMLA implemented the prediction models.
-JMLA conducted the experimental analyses, with NS supervising them, and HB providing critical revision for clinical intellectual content.
-JMLA produced the first draft, NS and HB revised it.
-All authors critically revised the content and approved the final version for publication.
+5.5 Ethics statement
+This study used publicly available and deidentified datasets.
+No direct patient interaction or intervention was involved.
+As these datasets are released under established data use agreements and have been ethically approved for secondary research, no additional ethics approval was sought.
+Appendix A Effect of the binary NaN mask on model performance
+For the ECG waveforms + clinical routine data setting (deep learning), our main results achieved 0.8256 (0.8222, 0.8288) for the diagnoses model and 0.9115 (0.8991, 0.9222) for the deterioration model when using a binary NaN mask column.
+In this section, we present additional experiments as a small ablation study to demonstrate that omitting the binary NaN mask column leads to inferior performance.
+Specifically, without the mask, the models achieved 0.8150 (0.8113, 0.8185) and 0.8986 (0.8864, 0.9094), respectively.
+Appendix B Advantages of multimodal integration in clinical settings
+In Table 5, we present a more finegrained performance comparison, which disentangles the overall performance according to ICD chapters in the diagnostic case and according to three categories, see Table 5 for the assignment of individual deterioration labels to these categories.
+In Table 6, we show a detailed breakdown of the model performance within the deterioration case.
+We focus on highlighting the added value of including full waveforms (ECG W.
+& CR) as opposed to waveform features (ECG F.
+& CR).
+We indicate the relative improvement of the former over the latter setting in the final column in both tables.
+Most notably, including full waveforms yields performance improvements for every single entry in both tables.
+Table 5: Detailed breakdown of model performance for different ICD-10 categories and clinical deterioration outcomes.
+Unimodal settings are CR: features derived from clinical routine data, ECG F.: ECG features, and ECG W.: ECG waveforms.
+Multimodal settings are ECG F.
+& CR: ECG features and clinical routine data, as well as ECG W.
+& CR: ECG waveforms and clinical routine data.
+Finally, Relative Improvement (%): relative improvement of ECG W.
+& CR against ECG F.
+& CR.
+CR ECG F. ECG W. ECG F. & CR ECG W. & CR Relative Improvement (%)
+IX: Circulatory 0.7645 0.7153 0.8050 0.8050 0.8761 8.83
+II: Neoplasms 0.8109 0.5886 0.8101 0.8101 0.8672 7.05
+XIX: Injury, Poisoning 0.7425 0.6262 0.7712 0.7712 0.8479 9.95
+I: Infections 0.7485 0.6258 0.7769 0.7769 0.8346 7.43
+IV: Endocrine 0.7723 0.6401 0.7920 0.7920 0.8336 5.25
+III: Blood 0.8048 0.5992 0.8143 0.8143 0.8335 2.36
+XI: Digestive 0.7510 0.5902 0.7604 0.7604 0.8270 8.76
+XIV: Genitourinary 0.7743 0.6053 0.7766 0.7766 0.8253 6.27
+X: Respiratory 0.7300 0.6581 0.7643 0.7643 0.8226 7.63
+V: Mental 0.7398 0.6456 0.7619 0.7619 0.8111 6.46
+XII: Skin 0.6951 0.6284 0.7153 0.7153 0.8087 13.06
+VI: Nervous 0.7088 0.5870 0.7183 0.7183 0.7955 10.75
+XVIII: Symptoms 0.6983 0.6090 0.7168 0.7168 0.7813 9.00
+XIII: Musculoskeletal 0.6786 0.5939 0.6930 0.6930 0.7405 6.85
+Clinical Deterioration 0.7729 0.7135 0.7515 0.8294 0.9070 9.36
+ICU Admission 0.8738 0.7377 0.7474 0.8865 0.9063 2.23
+Mortality 0.8792 0.7579 0.8160 0.8920 0.9168 2.78
+Table 6: Detailed breakdown of model performance category for all investigated sets of input features in the deterioration setting.
+Unimodal settings are CR: features derived from clinical routine data, ECG F.: ECG features, and ECG W.: ECG waveforms.
+Multimodal settings are ECG F.
+& CR: ECG features and clinical routine data, as well as ECG W.
+& CR: ECG waveforms and clinical routine data.
+Finally, Relative Improvement (%): relative improvement of ECG W.
+& CR against ECG F.
+& CR.
+We summarize the first six entries as ‘Clinical Deterioration’, the following two as ‘ICU Admission’ and the remaining seven as ‘Mortality’.
+CR ECG F. ECG W. ECG F. & CR ECG W. & CR Relative Improvement (%)
+Severe hypoxemia 0.4834 0.5311 0.7030 0.5055 0.6980 38.08
+ECMO 0.7873 0.6126 0.7768 0.8991 0.9355 4.05
+Vasopressors 0.8241 0.7759 0.7672 0.8551 0.9239 8.05
+Inotropes 0.6550 0.7900 0.6184 0.8064 0.9400 16.57
+Mechanical ventilation 0.9373 0.7262 0.7673 0.9418 0.9590 1.83
+IHCA 0.9502 0.8454 0.8760 0.9682 0.9859 1.83
+ICU 24 hours 0.8809 0.7390 0.7443 0.8928 0.9147 2.45
+ICU overall 0.8666 0.7363 0.7506 0.8801 0.8979 2.02
+In-hospital 0.8371 0.8753 0.8715 0.8736 0.9423 7.86
+24 hours 0.9304 0.7267 0.8409 0.9304 0.9600 3.18
+7 days 0.9176 0.7266 0.8176 0.9286 0.9429 1.54
+28 days 0.8858 0.7329 0.7859 0.8946 0.9115 1.89
+90 days 0.8651 0.7590 0.7926 0.8792 0.8952 1.82
+180 days 0.8605 0.7464 0.8036 0.8727 0.8894 1.91
+365 days 0.8579 0.7387 0.7999 0.8646 0.8768 1.41
+Appendix C Dataset details
+Table 7: Definitions of labels for the diagnoses and deterioration models, including 1428 discharge diagnoses based on ICD10-CM codes in the diagnoses model, and 15 labels covering clinical outcomes such as mortalities, ICU admission, vasopressors, inotropes, mechanical ventilation, in-hospital cardiac arrest (IHCA), extracorporeal membrane oxygenation (ECMO), and severe hypoxemia in the deterioration model.
+Name Labels Details
+Diagnoses model (1428 labels, based on [12])
+Diagnoses 1428 Discharge diagnoses ICD10-CM codes
+Deterioration model (15 labels, based on [23])
+Mortalities 6 1, 7, 30, 90, 180, 365 days and in-stay
+ICU Admission 2 24 hours and in-stay
+Vasopressors 1 Within 24 hours: epinephrine, norepinephrine, vasopressin, dobutamine, dopamine, or phenylephrine
+Inotropes 1 Within 24 hours: epinephrine, dobutamine, or dopamine
+Mechanical Ventilation 1 Same day or next day ICD9/ICD10 codes: 9670, 9671, 9672, 5A1935Z, 5A1945Z, and 5A1955Z
+In-Hospital Cardiac Arrest (IHCA) 1 Same day or next day ICD9/ICD10 codes: I469, 4275, I462, V1253, I468
+Extracorporeal Membrane Oxygenation (ECMO) 1 Same day or next day ICD9/ICD10 codes: 3961, 3965, 3966, 5A1221Z, 5A1522G, 5A1522H, 5A15223, 5A1522F, 5A15A2F, 5A15A2G, and 5A15A2H
+Severe Hypoxemia 1 Oxygen saturation $\leq$ 85% within 24 hours
+Table 7 contains the considered tasks by description, number of labels, and additional details such as target criteria definition.
+Apart from diagnoses, we use the rest of the tasks defined as patient deterioration based on previous work from the literature that defines physiological deterioration [23].
+Although we consider diverse scores such as SOFA significant, calculating these markers requires a more detailed definition of downstream task prediction, which goes beyond the scope of maintaining homogeneity across tasks in this study.
+The curation of the dataset starts from MIMIC-IV-ECG-ICD [12], from which we obtained the ICD10 diagnoses codes for the complete MIMIC-IV-ECG [20] dataset, therefore, based on ECG records with available ED stay_id, we select the ones that happened between the patient in time and the 1.5 hours window.
+Mortalities at finite horizons were computed based on the date of death from [20], and for the computation of mortality per stay, we include discharge date times from the admissions table in [21] MIMIC-IV to account for final discharge, all of the mortalities dates were previously validated from the source with a follow-up up to 1 year after the patient’s last recorded discharge.
+ICU admission in 24 hours and per stay was computed based on the ICU stays table from [21].
+Mechanical ventilation and ECMO were computed given the procedures table from MIMIC-IV-ED given the patient admission and procedure date.
+Cardiac arrest was computed from the diagnosis tables from MIMIC-IV-ED and MIMIC-IV, where we first validated if the patient was discharged with the diagnosis code within the ED or hospital otherwise.
+We clarify that although mechanical ventilation, ECMO, and cardiac arrest do not provide an exact time of events, we carefully approximate the times with dates given the same day or the day after.
+Vasopressors and inotropes were computed from the pyxis table from MIMIC-IV-ED given the exact time of the applied medication.
+Finally, severe hypoxemia was computed from the vital signs table from MIMIC-IV-ED.
+In the deterioration setting, to not discard samples where the events happened before the 1.5 hours for a single task, e.g.
+severe hypoxemia within 1.5 hours, we include a special token for the task in the label space which we mask during training and evaluation in order not to negatively affect training and evaluation for other labels due to exclusion of entire patients, see D for more details.
+Table 8: Summary demographic characteristics across samples including gender distribution, age distribution (median with standard deviation and quartiles), and ethnicity distribution in the study population, presented as percentages.
+Name Values
+Gender (%)
+Female 68289 (52.9)
+Male 60806 (47.1)
+Age (%)
+Median years (SD) 64 (28)
+Quantile 1: 18-49 32634 (25.27)
+Quantile 2: 50-64 34376 (26.62)
+Quantile 3: 65-77 31821 (24.64)
+Quantile 4: 78-101 30264 (23.44)
+Ethnicity (%)
+White 78310 (60.66)
+Black 27680 (21.44)
+Hispanic 9230 (7.14)
+Asian 4700 (3.64)
+Other 9175 (7.10)
+Table 8 provides detailed demographic information on the study population, including gender, age, and ethnicity.
+The gender distribution shows a slightly higher proportion of females (52.9%) compared to males (47.1%).
+The age distribution is divided into four quantiles, with a median age of 64 years and a standard deviation of 28 years.
+Regarding ethnicity, the majority of the population is White (60.66%), followed by Black (21.44%), Hispanic (7.14%), Asian (3.64%), and individuals of other ethnicities (7.10%).
+This diverse demographic distribution ensures a comprehensive evaluation of the model across various patient groups.
+These variables were extracted from the ed stays table from MIMIC-IV-ED.
+Table 9: Summary of statistics across samples for biometric measurements, vital parameters, and triage acuity levels, including minimum, maximum, median, interquartile range (IQR), and number of records for each variable.
+Name Unit Minimum Maximum Median IQR Records
+Biometrics
+BMI kg/m2 0 99.2 27.7 8.70 71234
+Weight kg 20.41 367.99 77.11 27.56 74026
+Height cm 64.99 387.09 167.64 15.23 42084
+Vital parameters
+Temperature C 13.3 44.72 36.6 0.55 77179
+Heart rate bpm 0.0 570.0 84.0 30.0 153249
+Respiratory rate bpm 0.0 169.0 18.0 4.0 150525
+Oxygen saturation % 0.0 100.0 99.0 3.0 146140
+SBP mmHg 0.0 274.0 130.0 35.0 151346
+DBP mmHg 0.0 486.0 73.0 23.0 151254
+Triage
+Acuity level 1 5 2 1 126071
+Table 9 provides details on the biometrics, vital parameters, and triage values considered in this work, including the name, unit, minimum value, maximum value, median value, and the number of records.
+Dataset preprocessing includes weight and height conversion from pounds and inches to kilograms and centimeters.
+During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence.
+The specific thresholds for outliers were as follows: temperature below and above 50 and 150 Fahrenheit degrees respectively.
+Heart rate above 700, respiration rate above 300, oxygen saturation below 0 and above 100, diastolic and systolic blood pressures above 500.
+BMI above 100, weight below 20 or above 400, and height below 60 and above 400.
+The temperature unit was converted from Fahrenheit to Celsius, whereas weight from pounds to kilograms, and height from inches to centimeters.
+Biometrics features were extracted from the online medical records table from MIMIC-IV from which we match the closest value of a patient stay within 30 days of admission, otherwise missing.
+Vital parameters and acuity were extracted from the vital signs and triage tables of MIMIC-IV-ED.
+Table 10: Summary of laboratory values across samples considered in this study, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records for each parameter.
+Name Unit Minimum Maximum Median IQR Fluid Records
+Abs. Basophil Count K/uL 0.0 9.2 0.04 0.03 Blood 36837
+Abs. Eosinophil Count K/uL 0.0 15.33 0.09 0.15 Blood 36838
+Abs. Lymphocyte Count K/uL 0.0 88.08 1.47 1.12 Blood 36909
+Alanine Aminotransf. IU/L 1.0 1999.0 22.0 22.0 Blood 29999
+Albumin g/dL 1.0 6.7 4.0 0.79 Blood 28324
+Alkaline Phosphatase IU/L 5.0 1985.0 85.0 51.0 Blood 29923
+Asparate Aminotransf. IU/L 1.0 1996.0 29.0 29.0 Blood 29995
+Bands % 0.0 61.0 0.0 3.0 Blood 6057
+Base Excess mEq/L -413.0 29.0 0.0 6.0 Blood 8011
+Basophils % 0.0 35.0 0.4 0.39 Blood 79386
+Bicarbonate mEq/L 2.0 50.0 24.0 5.0 Blood 79771
+Bilirubin, Direct mg/dL 0.0 45.5 1.0 2.5 Blood 1174
+Bilirubin, Total mg/dL 0.0 55.8 0.5 0.5 Blood 30098
+C-Reactive Protein mg/L 0.1 299.3 17.95 75.14 Blood 1108
+Calcium, Total mg/dL 0.0 79.2 9.2 0.79 Blood 30262
+Carboxyhemoglobin % 0.0 18.0 2.0 2.0 Blood 1078
+Chloride mEq/L 58.0 140.0 102.0 6.0 Blood 79855
+Creatine Kinase (CK) IU/L 8.0 1998.0 115.0 158.0 Blood 9588
+CK, MB Isoenzyme ng/mL 1.0 497.0 3.0 4.0 Blood 9272
+Creatinine mg/dL 0.0 43.0 1.0 0.5 Blood 82850
+Eosinophils % 0.0 75.0 1.3 2.1 Blood 79387
+Fibrinogen, Functional mg/dL 35.0 1348.0 298.5 161.5 Blood 2280
+Free Calcium mmol/L 0.48 3.31 1.08 0.12 Blood 1472
+Glucose mg/dL 3.8 1684.0 113.0 50.0 Blood 85022
+Hematocrit % 6.9 66.1 38.3 7.89 Blood 82490
+Hemoglobin g/dL 0.0 21.8 12.6 3.0 Blood 84890
+INR(PT) nan 0.75 24.0 1.1 0.30 Blood 44656
+Lactate mmol/L 0.3 25.2 1.8 1.20 Blood 32666
+Lymphocytes % 0.0 100.0 19.7 17.1 Blood 79387
+Magnesium mg/dL 0.2 45.0 2.0 0.40 Blood 30472
+Neutrophils % 0.0 99.0 69.9 19.59 Blood 79387
+Oxygen Saturation % 0.2 901.0 80.0 32.0 Blood 4838
+PT sec 8.2 150.0 12.0 3.5 Blood 44656
+PTT sec 16.6 150.0 30.6 6.99 Blood 44497
+Phosphate mg/dL 0.2 23.8 3.4 1.1 Blood 29888
+Platelet Count K/uL 5.0 1898.0 227.0 105.0 Blood 82260
+Potassium mEq/L 1.3 10.7 4.2 0.80 Blood 79857
+RDW % 10.6 33.8 13.8 2.19 Blood 82266
+Red Blood Cells m/uL 0.65 8.06 4.25 0.96 Blood 82266
+Sodium mEq/L 88.0 179.0 139.0 5.0 Blood 79865
+Troponin T ng/mL 0.0 17.36 0.06 0.13 Blood 41806
+Urea Nitrogen mg/dL 1.0 263.0 18.0 13.0 Blood 82522
+White Blood Cells K/uL 0.0 632.1 8.1 4.60 Blood 82367
+pCO2 mm Hg 8.0 189.0 44.0 17.0 Blood 8008
+pH units 5.0 9.0 6.0 1.0 Urine 20968
+Table 10 provides details on the laboratory values considered in this work, including the name, unit, minimum value, maximum value, median value, fluid type, and the number of records.
+During the data cleaning process, we removed outliers and introduced missing values at various thresholds to ensure model computational convergence.
+The specific thresholds for outliers were as follows: absolute basophil count above 20, absolute eosinophil count above 20, absolute lymphocyte count above 100, alanine aminotransferase above 2000, alkaline phosphatase above 2000, aspartate aminotransferase above 2000, creatine kinase above 2000, glucose above 2000, lactate above 2000, and platelet count above 2000.
+Laboratory values were extracted from the laboratory events table from MIMIC-IV, where for each record we matched all the patient records of the desired tests available between patient record in-time and the 1.5 hours window.
+Appendix D Models
+Table 11: S4 model hyperparameters
+Hyperparameter Value
+Block of layers 4
+S4 model copies 512
+S4 state size 8
+Optimizer Adam
+Learning rate 0.001
+Weight decay 0.001
+Learning rate schedule constant
+Batch size 64
+Epochs 20
+Table 11 outlines the hyperparameters employed in the S4 model.
+The architecture consists of four blocks of layers, with each block containing 512 copies of the S4 model.
+The state size within the S4 model is set to 8.
+For optimization, the AdamW optimizer is utilized with a learning rate and weight decay both set to 0.001.
+The learning rate schedule is maintained constant throughout training.
+A batch size of 64 samples is used for each training iteration, spanning a total of 20 epochs.
+The training objective is to minimize the binary cross-entropy loss.
+During training, we apply a model selection strategy on the best performance (AUROC) on the validation set which usually converges earlier than the final number of epochs.
+As previously described in C, with the goal of not discarding samples whose any of the deterioration targets happens before the 1.5-hour window or the target event cannot be clearly defined, we introduce a special token in the label space for which we account during and training and evaluation to be discarded.
+As of XGboost models, we fit these models with the default hyperparameter values in a binary setting for the later AUROC aggregation across settings, models which in principle do not intrinsically allow the inclusion of samples during training with the special token to be discarded.